### v2

# Mixtera: A Data Plane for Foundation Model Training

Abstract.
State-of-the-art large language and vision models are trained over trillions of tokens that are aggregated from a large variety of sources. As training data collections grow, manually managing the samples becomes time-consuming, tedious, and prone to errors. Yet recent research shows that the data mixture and the order in which samples are visited during training can significantly influence model accuracy. We build and present Mixtera, a data plane for foundation model training that enables users to declaratively express which data samples should be used in which proportion and in which order during training. Mixtera is a centralized, read-only layer that is deployed on top of existing training data collections and can be declaratively queried. It operates independently of the filesystem structure and supports mixtures across arbitrary properties (e.g., language, source dataset) as well as dynamic adjustment of the mixture based on model feedback. We experimentally evaluate Mixtera and show that our implementation does not bottleneck training and scales to 256 GH200 superchips. We demonstrate how Mixtera supports recent advancements in mixing strategies by implementing the proposed Adaptive Data Optimization (ADO) algorithm in the system and evaluating its performance impact. We also explore the role of mixtures for vision-language models.
PVLDB Reference Format:
PVLDB, 1(1): 1 - 1, 1.
doi:1
††This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 1, No. 1 ISSN 2150-8097.
doi:1
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/eth-easl/mixtera.
1. Introduction
Large language and vision models (often called foundation models) have become omnipresent in our daily life. They show enormous capabilities in a diverse set of tasks (Brown et al., 2020; Bommasani et al., 2022; OpenAI, 2024; Meta, 2024a; Radford et al., 2019), such as assistance with writing and coding, video understanding, and even agentic interaction with the world. The training of such language and vision models (LLMs/VLMs) presents new challenges for managing training data due to the ever-growing sizes of models and datasets. To achieve high accuracy, state-of-the-art models train over trillions of tokens from aggregated data collections such as RedPajama (Weber et al., 2024), Dolma (Soldaini et al., 2024), or FineWeb (Penedo et al., 2024a). For example, Meta’s Llama 3.3 70B model is trained on a corpus of 15 trillion tokens (Meta, 2024a, b). These collections are built based on data from various sources, such as Wikipedia or Common Crawl dumps.
Training data is typically stored on distributed filesystems in GPU clusters or data lakes in the cloud, and managed manually and ad hoc by ML engineers (Figure 2a). Selecting the right proportions out of this data with particular characteristics (e.g., language, topic, source) for training is critical for model performance (Chen et al., 2023; Ye et al., 2024; Xie et al., 2024). Individual users write ad hoc scripts to process the training data, filter relevant subsets with the properties of interest, often pre-tokenize it, and then mix it for their use case, e.g., they might want to train on 50 % data from Wikipedia and 50 % from movie subtitles. This reflects our experience working with ML researchers as part of the Swiss AI initiative which aims to develop open-source LLMs (ETH AI Center & EPFL AI Center, 2025). We confirmed through discussions with industry teams that such setups are common. This can get complex as training data may need to be mixed based on a variety of characteristics. For example, in addition to satisfying source data proportions (Wikipedia vs. movie subtitles in the previous example), we may also want the training data to be 80 % in English and 20 % in German. Recent works show that the data mixture should also be adjusted during training. The SmolLM2 model is trained with four stages of data mixtures (Allal et al., 2025). Algorithms such as Adaptive Data Optimization (ADO) (Jiang et al., 2024), Aioli (Chen et al., 2024a), PiKE (Li et al., 2025), and Skill-It (Chen et al., 2023) even propose to adjust the mixture dynamically based on the model behavior (e.g., loss per domain) during training. Figure 1 shows that using ADO can increase the performance of LLMs over a static mixture across different scales on the downstream HellaSwag benchmark (Zellers et al., 2019).
Today, there is no open-source system that automatically manages vast amounts of training data based on fine-grained characteristics. Implementing data filtering and subsequent mixing requires users to manually keep track of metadata. At least in the open-source world, this is typically done as part of the directory structure of the filesystem, e.g., there is one subdirectory per source dataset, and then we sample from each directory (Figure 2a). This approach is limited because each data sample has multiple properties that can be used to determine whether it should be used for training. Filesystems fundamentally do not offer the right interface for managing training data and mixing, as they do not provide declarative query interfaces or a native way to track which model was trained on what data. Running a data processing job to fully materialize the mixed training set for each training run has lots of overhead and leads to data duplication. A streaming-oriented approach that selects, tokenizes, and mixes data on the fly provides much more flexibility and eases experimentation. Utilizing a full-fledged DBMS for tracking data properties would burden ML engineers with complexities such as database administration, schema design, and performance tuning. We need a lightweight data plane that enables users to declaratively query and mix data based on arbitrary properties, independent of the filesystem structure, and to adjust this mixture dynamically.
| Mixtera | HF Datasets | WebDatasets | Mosaic Streaming | |||||||
| File formats |
webdataset |
| Static filtering | declarative | using map UDFs | using map UDFs | using map UDFs | ||||||
| Static mixtures | on all properties | on filesystem dirs. | on filesystem dirs. | on filesystem dirs. | ||||||
| Dynamic mixtures | on all properties | no | no | no | ||||||
| Native 3D parallelism | yes | yes, manual rank handling | data parallel only | yes, for specific rank order | ||||||
| Checkpointing | yes | using TorchData | by replaying, not natively | yes |
We present Mixtera, a data plane that can be deployed on top of existing LLM/VLM training data collections. It is a centralized, read-only layer and can be declaratively queried from training clients. As shown in Figure 2b, it reduces the amount of materialization and data scripting necessary for training jobs. While existing data loaders as outlined in Table 1 are limited by filesystem constraints, Mixtera supports arbitrary properties independent of the filesystem, and makes it easy for model engineers to experiment with different filter criteria, fixed learning curricula, and fully dynamic mixing algorithms that learn the mixture during training. Mixtera follows a client-server model. For a training job, the server, which indexes and manages the samples, first statically filters out the relevant samples, and then continuously distributes chunks to the clients. Chunks are fixed-size collections of pointers to samples in files, and adhere to the current mixture. The clients then fetch the relevant samples and return the relevant data to the training loop. We design and implement Mixtera tailored to the needs for foundation model training, and make the following key contributions:
Mixtera enables users to declaratively specify mixtures across properties independent of the filesystem structure and training frameworks by indexing all samples and their properties.
Mixtera enables dynamically changing the data mixture and the properties determining the mixture during training. It achieves this by generating and streaming chunks, i.e., fixed-size lists of pointers to samples following the current mixture.
We show Mixtera does not bottleneck training and is versatile, enabling model accuracy improvements for text-only and text-vision models. As an example, we demonstrate how to implement the ADO dynamic data mixing algorithm in Mixtera and its positive impact on model accuracy.
2. Background
Foundation models are large-scale deep learning models suitable for a variety of tasks (Bommasani et al., 2022; Competition and Markets Authority, 2013). We focus on text-generation models, i.e., autoregressive large language models (LLMs) and multimodal vision-language models (VLMs). As of 2025, most such models are based on the Transformer architecture (Vaswani et al., 2017). They are trained on huge corpora of training data in a self-supervised manner to maximize the likelihood of predicting the tokens of a training sequence.
Training phases. Training is structured into pre-training and post-training phases. In pre-training, we train a randomly initialized model on a general purpose data corpus (Section 2.1) to derive a base model. In post-training, common steps include supervised finetuning (SFT) and alignment. The data used in SFT is human-selected for a specific task. In this paper, we focus on the pre-training use case.
Distributed training. Training foundation models requires distributing computation across multiple GPUs. Training frameworks typically employ 3D parallelism (Ben-Nun and Hoefler, 2019; Hoefler et al., 2024), consisting of pipeline parallelism (PP), i.e., partioning the model layers between devices (Narayanan et al., 2021; Huang et al., 2019; Narayanan et al., 2019), tensor parallelism (TP), i.e., splitting individual tensor operations within layers across devices (Shazeer et al., 2018; Shoeybi et al., 2019), and data parallelism (DP), i.e., replication of the model across device groups. PP and TP together are referred to as model parallelism. Nodes within the same DP group process identical inputs, while nodes across DP groups receive different data. As an extension to DP, fully-sharded data parallelism (FSDP) shards model parameters, gradients, and optimizer states across data-parallel workers (Zhao et al., 2023).
2.1. Training Data and Data Mixing
The data used for pre-training stems from data collections that include data from various sources, such as Wikipedia, Common Crawl dumps, or arXiv papers. Public examples of such collections include RedPajama (Weber et al., 2024), Dolma (Soldaini et al., 2024), and FineWeb (Penedo et al., 2024a). Besides the aggregation of data from different sources, data engineers typically clean the data, which typically involves deduplicating, filtering (e.g., the removal of personal identifiable information - PII), and applying classifiers to the data samples (e.g., to obtain a toxicity score for each sample). Longpre et al. (Longpre et al., 2024) and Penedo et al. (Penedo et al., 2024a) provide a great overview of the processing steps and their impact.
Data properties and mixtures. Each sample in the resulting collection has properties, such as its source (e.g., Wikipedia) or its language (e.g., English). Beyond the filtering operations, ML engineers need to define a data mixture. A mixture describes how the data is mixed based on their characteristics, i.e., we might train on 50 % data from Common Crawl and 50 % from movie subtitles. The data can be combined based on multiple characteristics simultaneously. For instance, besides Common Crawl and movie subtitles, we might also use 80 % French and 20 % Italian data. The granularity on which the mixture is ensured depends on each training setup and there is no common agreement. For example, a mixture could be ensured within a batch, or across a window of several batches.
Mixing algorithms. Selecting the best mixture is critical for model performance (Chen et al., 2023; Ye et al., 2024; Xie et al., 2024; Shen et al., 2024). We differentiate static mixtures, i.e., mixtures that remain constant over the entire training job, and dynamic mixtures, i.e., mixtures that change during the training job. In the last years, several algorithms for finding the best static mixture or how to adjust the mixture dynamically during training have been proposed. Algorithms such as DoReMi (Xie et al., 2024) or the data mixing laws (Ye et al., 2024) find the a static mixture via small proxy models.
Curriculum learning is an example of a pre-defined dynamic mixture. Xu et al. (Xu et al., 2024b) order samples from easy to hard to improve alignment. The SmolLM2 model was trained on 4 stages of mixtures (Allal et al., 2025). Multilingual models are often first trained on English data, and then data from other languages is included (Richburg and Carpuat, 2024; Xu et al., 2024a).
Beyond such pre-defined schedules, there is also work on adapting the data mixture to the model training dynamics, e.g., by increasing the weight of data domains which have high loss. Albalak et al. (Albalak et al., 2023) use a multi-armed bandit strategy to adjust the mixture. Skill-it orders “skills” based on model feedback (Chen et al., 2023). Aioli builds upon Skill-it and provides a unified framework for estimating the best mixture during training (Chen et al., 2024a). PiKE relies on gradient interactions (Li et al., 2025). In this paper, we use Adaptive Data Optimization (ADO) (Jiang et al., 2024) as an example of a dynamic mixing algorithms.
2.1.1. Adaptive Data Optimization
Adaptive Data Optimization (ADO) is a dynamic mixing algorithm that adjusts the data mixture during training based on the model’s learning progress on each domain (Jiang et al., 2024). The key idea is to prioritize domains where the model shows rapid improvement while considering how much each domain contributes to its own progress. ADO uses neural scaling laws to model how the loss of each domain decreases with the number of training samples . To this end, it fits a power law for each domain. Here, represents the irreducible loss of the domain, is a scaling factor, and determines how quickly the loss decreases. The parameters are re-fitted during training. The algorithm combines two components to determine the mixture weights. First, it estimates the learning speed for each domain using the derivative of the scaling law. Second, ADO maintains a credit assignment score that indicates how much each domain contributes to its own progress, based on its recent sampling frequency. These components are combined with a prior (initial) distribution to compute an intermediate preference distribution . To ensure stability, the final distribution is then computed as a weighted average between and ’s temporal average. Additionally, ADO enforces a minimum sampling probability for each domain. For more details, we refer to Jiang et al. (Jiang et al., 2024).
3. Current Challenges
We identify three challenges in the status quo of training data management with current open-source infrastructure.
Challenge 1: High engineering effort for data preparation. The current approach to preparing training data sets involves many manual offline steps with general-purpose data processing and scripting frameworks (Figure 2a). For the offline cleaning step (Section 2.1), ML engineers typically leverage data processing frameworks like Spark (Zaharia et al., 2016), Beam (Akidau et al., 2015), Data-Juicer (Chen et al., 2024b), or datatrove (Penedo et al., 2024b).
The subsequent data mixing can happen offline or online. If data is mixed offline, engineers write ad hoc scripts which create a new mixed copy of the cleaned dataset for each training run. This makes it difficult to get a quick sense of how a data mixture will impact model training when exploring different mixing policies. Some existing data loaders support online data mixing. However, they only implement it across directories, i.e., they assign a weight to each directory and sample data according to the weights. Hence, the directories need to reflect the property we want to mix on, e.g., one directory per language or data source. This inflexible approach neither supports switching the property we mix on nor supports specifying hierarchical mixtures across arbitrary properties, e.g., specify a proportion between Wikipedia and movie subtitles, and between English and German (c.f. Table 1).
Challenge 2: Inefficient data management on filesystems. Training data is typically stored and managed as files without a proper data management system, which can lead to storage overhead, performance bottlenecks, and consistency issues. Both the source data and mixed training set are typically stored in formats such as jsonl or parquet files in a shared, distributed filesystem. Existing data loaders just wrap around the filesystem and are therefore limited by its constraints (Table 1). Filesystems are commonly used because current database management systems are not natively built for foundation model training data (Wang et al., 2023). Using a proper DBMS for such data would require ML developers to define table schemas and to orchestrate dataflow from the DBMS to the model training, in addition to administration overhead to operate the DBMS itself. This situation is problematic, as (i) both the metadata and actual payloads are commonly duplicated across training jobs, (ii) filesystems do not provide an efficient interfaces to query data, (iii) there is no native way of tracking which model was trained on which data if it is accessed via general-purpose filesystem calls. Some frameworks like Megatron (Shoeybi et al., 2019; Narayanan et al., 2021) even require pre-tokenized data, which leads to duplication of the even bigger tokenized data files.
Challenge 3: Rapidly evolving research. How to train the best model on a given dataset is an active area of research, with many new techniques such as dynamic mixing emerging. Offline preparation of the mixture or online mixing based on fixed directory weights does not support dynamic mixture at all. Even for researchers who are familiar with the latest mixing techniques, implementing modern mixing algorithms in a training pipeline is a painful, tedious, and error-prone task. The codebases for mixing algorithms are often tailored directly to the training framework, as well as the data collection and properties used in the respective papers. This hinders the adoption of new methods and makes researching, reproducing, and comparing different strategies difficult.
4. Mixtera’s Design
We propose to address these challenges for training data management by building Mixtera, a foundation model training data plane. We derive the following design goals for such a system.
Goal G1: The system should implement a centralized data layer that users can conveniently and declaratively query to mix data across arbitrary properties, independent of the filesystem structure.
Goal G2: The system needs to be lightweight, i.e., not require many components to set up and be easily integratable into existing training setups.
Goal G3: While being user-friendly and flexible, the system needs to ensure high-throughput and determinism (reproducibility).
Goal G4: The system must support adjusting the mixture dynamically during the training.
4.1. System Overview and Design
TODO!
Inspired by the Lakehouse architecture (Armbrust et al., 2021), we design Mixtera as a read-only layer that can be deployed on top of existing training data collections, which are typically stored in a distributed filesystem. Figure 3 shows the architecture of the system. Mixtera manages a centralized database of metadata (i.e., properties such as language or source dataset) of all training data samples (G1). A sample can be a piece of text (for LLM training) or a text-image pair (for VLM training). Mixtera assigns every sample a unique ID and properties instead of treating the training data as a blob of homogeneous, contiguous data. It allows users to declaratively query the relevant samples for a training. To be lightweight (G2), Mixtera does not reorganize or modify the data files on disk. For model training, it provides a standard iterator that can be used, e.g., in conjunction with a torch.DataLoader. Mixtera is agnostic to the model training framework, supports training interruptions using checkpoints, and ensures determinism (G3) through careful shuffling, i.e., for identical queries, Mixtera always provides data in an identical order, which is important for reproducibility as well as for debugging issues like loss spikes (Zhuang et al., 2022; Cooper et al., 2022; Qian et al., 2021; Karamcheti et al., 2021; Team, 2025). It supports adjusting the mixture during training (G4) by transferring chunks (lists of pointers to samples) whose mixture can change over time.
[ frame=lines, framesep=2mm, baselinestretch=0.8, bgcolor=LightGray, fontsize=, linenos ] python client = MixteraClient("127.0.0.1", 8080) job_id = "test_job" query = Query.for_job(job_id).select(("license","==","CC")) mixture = StaticMixture( MixtureKey("language": ["JavaScript"]): 0.7, MixtureKey("language": ["HTML"]): 0.3 , chunk_size=1024) qea = QueryExecutionArgs(mixture=mixture, num_workers=4, dp_groups=1, nodes_per_group=1) rsa = ResultStreamingArgs(node_id=0, dp_group_id=0, job_id=job_id) ds = MixteraTorchDataset(client, query, rea, rsa) dl = torch.utils.data.DataLoader(ds, batch_size=1024, num_workers=4)
[…]
Query interface. In Figure 4, we show an example of a query that statically selects only Creative Commons data, and then mixes HTML and JavaScript data in a 70:30 ratio during training. Mixtera’s implementation takes care of executing the query and obtaining the samples without needing to worry about correctness, even in distributed training. The user only needs to provide the ID of the node and its data parallel group, which is obtained from the training framework. Mixtera currently allows to express static filter operations on properties, and static as well as dynamic mixtures across all properties (G4). For more details on supported mixture types, we refer to Section 5.2.2.
Dataflow. Mixtera follows a server-client model. As shown in Figure 3, the server runs on a node and each training node runs client instances. Users first register samples at the server to populate the metadata database. They then can submit queries. A query is executed at the server in two phases. First, Mixtera applies static filters from the query (e.g., English-only) to obtain all samples eligible for training. Second, during training, the server distributes chunks of that query result to the client(s), which specify which samples to train on. The server ensures that the chunks are distributed correctly, i.e., tensor- and pipeline parallel stages receive the same input data. The server generates chunks according to the current mixture, i.e., it iteratively takes samples from the query result generated in the first phase of query execution, such that data in the chunk follow the current mixture. As an iterable data loader, Mixtera faces the challenges of determinism and checkpointing. We address this by shuffling based on the query and support to load/store the query state.
Chunks. Mixtera does not store the sample payloads, but rather only the metadata of each sample. During training, the nodes receive chunks. A chunk is a fixed-size collection of pointers to samples in files. They tell the client which samples which file to load and train on (e.g., sample 10 in file wikipedia.jsonl.zst). The files can be local, in a distributed filesystem, or cloud storage.
Storing and distributing pointers to samples instead has several advantages. First, users can store data in their locations of choice (e.g., an object store, or an distributed filesystem). Chunks are independent of the filesystem structure (G1). Second, it allows Mixtera to support dynamic mixtures (G4), as the data composition of chunks can change over time. Third, the pointer-model avoids creating a data fetching bottleneck at the Mixtera server. The server only creates chunks and each client fetches the data they need. Fourth, we avoid a lock-in effect on Mixtera and allow for easy adoption on existing data collections (G2). Last, we natively support other modalities like images, which would not be straightforward to ingest at scale into a database.
Mixtures. Users can use any subset of sample properties to define a mixture. Users can even change the properties that they mix on during training with dynamic mixture schedules. To handle this, Mixtera implements a MixtureKey abstraction that allows users to describe which samples they want to use. The keys also implement flexible property matching to define which samples to use during chunk generation (Section 5.2.2).
High-throughput data fetching. The challenge of using chunks is that we need to handle suboptimal data layout. Files may have arbitrary property distributions, or might follow a partial structure (e.g., a file only contains Wikipedia data, but the languages are distributed randomly). We cannot re-organize data within the existing files. Formats like jsonl were not built with random access in mind, but chunks force clients to load individual samples from files. To avoid data stalls (G3), we implement chunk generation to use subsequent sample ranges in files if the samples have the same properties and engineer Mixtera’s implementation to fetch those ranges as efficiently as possible (Section 5.3).
Open source. Mixtera comes as a Python package that provides the entrypoint for the server and abstractions for the client. Its codebase, consisting of approximately 11 k lines of Python and C++ (excluding tests), is open-source111Available at https://github.com/eth-easl/mixtera.. It is rigorously tested with a full set of unit and integration tests. We are continuously working on enhancing the system and welcome contributions.
5. Implementation
We explain how sample metadata is managed (Section 5.1) at the server, how it executes queries and creates chunks (Section 5.2), how those chunks are parsed at the client (Section 5.3), and give details on Mixtera’s integration into training frameworks (Section 5.4).
5.1. Sample Management
Mixtera manages samples in a MixteraDataCollection (MDC). It keeps track of all samples and their properties. The MDC leverages DuckDB (Raasveldt and Mühleisen, 2019) as the robust and flexible metadata management system underlying it. Every registered sample has properties according to a schema defined by a MetadataParser, e.g., language. Representing the structure of training data, Mixtera initializes the database with three tables for source datasets, files, and samples. The source dataset can also be ignored and stored as a schema property instead, to support collections such as The Pile (Gao et al., 2020) where data from different datasets is contained within the same files.
MetadataParsers. A MetadataParser is a class that defines a schema for a data collection and, given a sample, extracts and returns the properties according to the schema. A schema is a list of properties which have a type (e.g., string or enum), a nullable field, and a multiple field, describing whether a single sample can take multiple values for this properties, e.g., multiple languages. Mixtera maps this Python schema to a proper database schema, comes with a set of pre-defined parsers for common datasets, and enables users to define custom parsers. For example, there is a parser for The Pile (Gao et al., 2020). Every sample has the property pile_set_name describing the source dataset. This is a non-nullable, non-multiple, enum property, since we know the limited set of possible values.
Data types. Based on the parser, Mixtera adjusts the DuckDB sample table. Properties marked as multiple get mapped to DuckDB’s list type, e.g., a multiple string property gets mapped to VARCHAR[]. Enum properties are appropriately inserted into DuckDB. Using enum instead of strings for properties where all values are known beforehand speeds up query execution in DuckDB.
Insertion. When inserting (registering) data, the MDC accumulates all relevant files, and prepares the DuckDB table schema. Then, all files are inserted into the files table, and a pool of workers in parallel processes all files using the MetadataParser to extract the sample metadata. We then aggregrate the worker results and insert it into the database, as DuckDB does not support insertion from multiple Python process workers. We found that converting the worker results to columnar pyarrow in-memory tables and then inserting into the samples table has the highest throughput.
5.2. Server-Side Query Execution
After receiving a query from the client, the server executes it in two phases. The first phase is performed via the MDC DuckDB-abstraction and applies static filters to identify all potentially relevant samples, e.g., all non-English samples are filtered out, and groups consecutive samples into intervals. The second phase constructs a specialized data structure called ChunkerIndex that enables efficient, mixture-aware chunk generation.
5.2.1. SQL generation and interval detection
After receiving an object representation of the query (c.f. Figure 4), similar to ORM frameworks like sqlalchemy, the Mixtera server generates a base SQL query from this object. This query returns a table in which each row represents a sample that the user is interested in. Mixtera ensures that the generated SQL matches the MDC’s table schema, e.g., whether a property can have multiple values or not.
A key challenge for Mixtera is efficient random access to samples within files. File formats like jsonl or parquet are optimized for sequential reading rather than random access. To address this, Mixtera implements an interval-based approach: the server wraps the base filtering query in an outer query that identifies continuous ranges of samples sharing identical properties within the same file. Consider the following example result of a base filtering query:
| Sample ID | File ID | Language | License |
| 1 | 1 | JavaScript | MIT |
| 2 | 1 | JavaScript | MIT |
| 3 | 1 | JavaScript | MIT |
| 4 | 1 | Python | Apache |
| 5 | 1 | Python | Apache |
| 1 | 2 | Python | Apache |
Instead of treating these as six individual samples, Mixtera identifies three intervals:
Interval 1: Samples 1-3 (File 1, JavaScript, MIT)
Interval 2: Samples 4-5 (File 1, Python, Apache)
Interval 3: Sample 1 (File 2, Python, Apache)
Even though samples 4-5 (file 1) and 1 (file 2) share the same properties (Python, Apache), they are in different files and thus form separate intervals. The primary key is formed by the sample and file ID together. Mixtera constructs a SQL query that processes the data in multiple stages:
First, Mixtera establishes a Common Table Expression (CTE) named base_data that contains the filtered samples: {minted}[frame=lines,framesep=2mm,baselinestretch=0.8,bgcolor=LightGray,fontsize=]sql WITH base_data AS ( – Our generated base filtering query here, e.g., SELECT * FROM samples WHERE license = ’MIT’),
Next, Mixtera identifies breaks in the sample sequence using window functions. The grouped_samples CTE calculates the difference between consecutive sample IDs within groups sharing the same properties: {minted}[frame=lines,framesep=2mm,baselinestretch=0.8,bgcolor=LightGray,fontsize=]sql grouped_samples AS ( SELECT *, sample_id - LAG(sample_id, 1, sample_id) OVER (PARTITION BY file_id, lang, license ORDER BY sample_id) AS diff FROM base_data), Here, a diff value of 1 indicates consecutive samples, while any other value indicates a break in the sequence.
The intervals CTE then groups the sequences into intervals: {minted}[frame=lines,framesep=2mm,baselinestretch=0.8,bgcolor=LightGray,fontsize=]sql intervals AS ( SELECT file_id, lang, license, SUM(CASE WHEN diff != 1 THEN 1 ELSE 0 END) OVER (PARTITION BY file_id, lang, license ORDER BY sample_id) AS group_id, MIN(sample_id) as int_strt, MAX(sample_id)+1 as int_end FROM grouped_samples GROUP BY file_id, lang, license, diff, sample_id) The group_id is incremented when there is a break in the sequence, creating unique identifiers for each interval.
Finally, Mixtera aggregates the results to get the final intervals: {minted}[frame=lines,framesep=2mm,baselinestretch=0.8,bgcolor=LightGray,fontsize=]sql SELECT file_id, lang, license, group_id, MIN(int_strt) as interval_start, MAX(int_end) as interval_end FROM intervals GROUP BY file_id, lang, license, group_id ORDER BY file_id, interval_start;
While we initially implemented the interval aggregation within Python, letting DuckDB optimize and parallelize the calculation of the interval is more efficient. This optimization can only improve I/O if samples within files are clustered by properties and not randomly distributed. For example, if a user has a file structure based on the source dataset and mixes based on that, Mixtera can heavily leverage the range optimization and just read, e.g., all Wikipedia samples from the sample file. If the user decides to now mix on language as well, there might only be local clusters within the files. Since Mixtera is read-only by design, it cannot re-shuffle data.
5.2.2. Chunk generation
The chunk generation algorithm is based on the ChunkerIndex data structure, which organizes sample ranges by their properties. This data structure is generated based on the output from the interval query. Before explaining its construction, we first introduce the key abstraction that enables flexible property matching: MixtureKey.
MixtureKey abstraction. A MixtureKey represents a set of properties and their values, e.g., language being English and license being MIT. A property in a key can have multiple values. A key matches another if their lists of values intersect for all properties they share. This matching is crucial because the resulting interval table from DuckDB contains the full cross-product of all properties—a sample might have values for language, license, size, topic, and more—while a mixture specification typically considers only a subset of these properties. For example, a mixture might specify language:English to select English text regardless of license, while the underlying samples are stored with full property information (e.g., language:English,German;license:CC and language:English;license:MIT). The flexible matching enables finding all samples that satisfy certain properties while ignoring irrelevant attributes they might have. It also allows us to define mixtures on multiple properties with multiple values, instead of being limited to a single property (c.f. Section 3). To ensure deterministic behavior, we implement a total ordering over keys based on the number of properties, property names, and their values. We sometimes refer to a specific MixtureKey as a domain, e.g., the key for language:English defines the domain for English samples.
The ChunkerIndex. The ChunkerIndex is a nested mapping from MixtureKeys to sample locations. We refer to keys in the ChunkerIndex as component keys. For each unique combination of all available properties and values, the index maps to dataset IDs, which in turn map to file IDs and finally to sorted lists of intervals of samples which share the properties and values as specified in the MixtureKey. While the index maintains the complete property information of samples, the matching semantics allows us to efficiently query arbitrary subsets of these properties. Consider a simplified example where we represent MixtureKeys as strings. A fragment of the ChunkerIndex might look like: {minted}[frame=lines,framesep=2mm,baselinestretch=0.8,bgcolor=LightGray,fontsize=]python "language:JavaScript,HTML;license:MIT": ds_1: file_1: [(1,4), (10,15)], # half (right) open ranges file_2: [(1,2)] , "language:Python;license:Apache": ds_1: file_1: [(1,2)]
In this example, a query for language:JavaScript would match the first key despite it having the additional license property and two assigned languages. This demonstrates how the MixtureKey matching allow to work with the full property/value cross-product in the index while supporting mixtures on subsets of properties.
Building the ChunkerIndex. The index is built in parallel in a C++ extension, processing the interval table from DuckDB provided in Apache Arrow format. Operating on the Arrow table in Python would be too slow due to Global Interpreter Lock (GIL) constraints. Using multiprocessing to circumvent the GIL would require expensive pickling of nested dictionaries. Our C++ implementation uses multithreading and only acquires the GIL at the end when creating the final Python result object.
Each C++ worker thread maintains a local index for a subset of the data. For each row (interval), each worker constructs a C++-representation of the MixtureKey, inserts the interval into its local index under this key, maintaining sorted order within each file’s interval list. After parallel processing, the local indices are merged, combining interval lists while preserving their sorted order. In the end, we convert the index to Python objects, which often is the most expensive operation of this process.
Chunk generation. A mixture in Mixtera is represented as a mapping from MixtureKeys to their target proportions within a chunk. Given a chunk size that describes how many samples each chunk should contain, these proportions are converted to absolute counts following the largest remainders quota method. Users can set a mixture to be strict, requiring exact proportions, or best-effort, allowing to continue to generate chunks even if the mixture cannot be exactly fulfilled. The chunk generation algorithm (Algorithm 1) iteratively constructs chunks that conform to the current mixture. This algorithm supports dynamic mixture, as the mixture can be changed between chunks.
Similar to the ChunkerIndex, a chunk is also structured as a hierarchical dict with keys and files pointing to lists of ranges. This chunk can be parsed on a training node by loading the sample ranges from the files (Section 5.3). While the ChunkerIndex uses all properties/values as MixtureKeys (component keys), a chunk’s keys are identical to the mixture’s keys, which might only contain a subset of properties.
For each key in the mixture, the algorithm keeps track of how many samples we still need to put into the chunk that is currently being generated (remaining_sizes). For each key in the mixture (line 6), it checks whether it matches a component key in the chunker index (lines 7-8). If we find a match, we try and obtain samples (ranges) from the ChunkerIndex for this component key (lines 9+).
A call to obtain samples for a component key can return fewer samples than requested, e.g., if we are looking for JavaScript data and we need 5 samples, but we only have 3 JavaScript/MIT licensed samples, the according component key can only return 3 samples. Requesting samples is implemented as requesting ranges (from potentially multiple files) such that the overall number of samples in the returned list of ranges is . The lists of ranges per file are merged into the existing sorted list of ranges. Mixtera’s implementation uses Python generators that yield ranges of samples and accept the number of samples needed as input through the Python generator’s send mechanism. This hides the complexity of the take samples operation. It also allows for efficient, stateful iteration over the available ranges while maintaining control over how many samples are returned in each request. There is one generator per component key that returns the ranges containing samples based on the ChunkerIndex but ensures ranges are split such that never too much data is returned.
If after traversing all component keys we did not find sufficient samples for a key in the mixture, in strict mode, chunk generation fails. In best-effort mode, the algorithm redistributes any unfulfilled counts to the remaining mixture components proportionally to their original ratios. For example, if we need 100 JavaScript samples but only find 80, the remaining 20 samples would be proportionally distributed among other components. To avoid infinite loops, we only distribute samples to keys on which we were able to find any samples in the last iteration.
Mixtera’s implementation of this algorithm ensures determinism because (1) the mixture’s keys are processed in a consistent order and (2) when multiple component keys match a mixture key, they are considered in a deterministic order based on a seeded shuffle of all possible keys. This approach ensures that identical queries with identical mixtures always produce identical chunks, important for training debugging and reproducibility (Zhuang et al., 2022; Cooper et al., 2022; Qian et al., 2021; Karamcheti et al., 2021; Team, 2025).
Mixtera also supports an alternative, arbitrary chunk generation algorithm. It does not require any specified mixture and builds up chunks by iterating over the component keys, moving on to the next one when one key is depleted. This can be useful if the data is completely property-free and guarantees maximum throughput as all ranges are consecutive.
Mixture types. All mixture classes implemented in Mixtera share a common interface that converts their specifications into a mapping from MixtureKeys to sample counts per chunk, used by the chunk generation algorithm:
Static Mixture: Users explicitly specify fixed proportions for different property combinations (Figure 4). This supports arbitrary properties and is not limited by, e.g., directory boundaries.
Inferring Mixture: Automatically derives mixture proportions from the data distribution in the query result: This is useful when users want to maintain the natural distribution of properties.
Hierarchical Mixture: An advanced static mixture that allows to specify nested property relationships. For example, users can define that 50 % of the data should be legal texts, and within that, 60 % should be in English and 40% in French. Mixtera automatically flattens this hierarchy into appropriate MixtureKeys.
Mixture Schedule: A “meta mixture” that allows for temporal changes in mixture composition by defining a sequence of mixtures that activate at specific training steps. This enables curriculum learning with predefined schedules.
Dynamic Mixture: Allows adaptation of mixture proportions during training based on feedback (e.g., loss) from the model. If an algorithm is already supported by Mixtera (e.g., ADO), it can be used directly.
Chunk distribution. In distributed training, it is important to guarantee that all nodes within the same data parallel group operate on exactly the same input tensors. Mixtera’s ChunkDistributor wraps around the chunk generation component, and hands out chunks correctly to the training nodes, i.e., the same chunks in the same order to nodes within the same group, and different chunks to nodes in different groups for data parallelism. To this end, the clients need to register at the server with their respective node and group identifiers. To avoid redundant serialization overhead, the distributor caches chunks in serialized form until all nodes in a group have received them.
5.2.3. Networking
We implement a TCP-based client-server protocol. The server uses Python’s asyncio framework to handle multiple concurrent client connections. The protocol is message-based, with each message consisting of a task identifier followed by task-specific payload data. Tasks for example include the execution of a query or sending out a new chunk to a client. To handle network issues gracefully, the client implementation includes automatic reconnection with exponential backoff and configurable timeouts. Larger data transfers, such as chunk transmission, use efficient streaming to avoid memory issues.
5.3. Client-Side Reading
Mixtera’s client-side abstractions offer a generator that yields data samples for a query at the server. This generator internally follows a two-level nested iteration pattern: an outer iteration over chunks and an inner iteration over samples within each chunk. The outer iteration hides the complexity of network transfer and chunk generation, while the inner iteration hides the complexity of going from pointers in the chunk to actual samples. We discuss this inner step, i.e., how, given a chunk, we yield sample payloads.
Processing modes. A chunk can be processed in three mixture processing modes with different trade-offs. The modes influence in what order samples are yielded, i.e., in what granularity the mixture is guaranteed, and whether string samples or tokenized sequences are yielded. All modes begin by instantiating one active iterator per property combination. These iterators traverse files and ranges for their respective properties in a randomized order while maintaining sequential reading within consecutive ranges for I/O efficiency.
Overall mixture mode. This mode processes active iterators in a randomized round-robin fashion until depletion. This ensures the mixture ratio is maintained at the chunk level.
Window mixture mode. This mode guarantees the mixture on a window smaller than the chunk size. Similarly to chunk generation, we determine how many samples per property we yield within a window. We then go through the properties in a randomized, round robin fashion until a window has been yielded, and start again. This mode can operate in best-effort (continues after mixture cannot be guaranteed) or strict mode (stops at the first window where the mixture cannot be maintained). In strict mode, the number of overall samples yielded from the chunk might be smaller than the chunk size.
Tokenized mixture mode. This mode extends the strict window mode. It wraps the active iterators with a tokenizing iterator that takes the incoming string samples, tokenizes them, and yields tokenized samples (integer lists) with the correct sequence length. By setting the window size equal to the chunk size we guarantee that each chunk at least yields one window of tokenized samples. This mode is important for handling imbalanced collection with significantly varying text lengths. For example, in the The Pile (Gao et al., 2020), an average sample from the Books3 domain is much longer than one from the PubMed Abstracts domain.
Without tokenization, one string sample may correspond to potentially order-of-magnitudes more tokenized samples, which can heavily disturb the actual mixture seen during training. While this mode ensures precise mixture ratios at the token level, it may result in partial utilization of longer samples. This is an inherent issue of unbalanced datasets, and while Mixtera provides flexibility to handle it, the best approach is to process the datasets such that samples are (roughly) of similar size.
File reading. The active iterators wrapped by the processing iterators shuffle the file order but maintain sequential reading within files. This acknowledges that formats like jsonl and parquet are optimized for sequential rather than random access, and enables us to linearly iterate through the sorted lists of ranges per file. The complexity of reading different file formats internally is hidden by abstractions for each format. Using the xopen library we support both compressed and uncompressed jsonl. We optimize the reading of parquet files by calculating and loading only the relevant the row groups, and build upon pyarrow’s parquet-batched-reading implementation. WebDatasets is the only format supporting random access to samples, and we implement support using the wids library. The format also gives us the option to store text-image pairs, as it can contain different modalities. To mitigate latency from initial file operations that we observed on distributed filesystems, Mixtera employs a prefetching iterator that uses background threads to hide file opening latency.
Determinism. All random operations are seeded based on the current chunk, ensuring identical behavior across nodes. Combined with the server-side chunk generation and distribution, this guarantees that all clients within a data parallel group yield exactly the same samples in exactly the same order. This property is crucial for both reproducibility across runs and correctness in distributed training. We validated this using a suite of integration tests as well as the dataloader verification test provided by nanotron.
Dataset abstractions. Training frameworks typically require specific dataset interfaces that supports multiprocessing with worker processes. Besides a general purpose interface, Mixtera offers a class extending torch’s IterableDataset, and a class compatible with the huggingface API. Each worker process at each node operates on its own chunk.
5.4. Framework Integration
Mixtera integrates into the training framework for checkpointing and transferring model feedback (e.g., per-domain loss).
Checkpointing. Mixtera’s API offers a function to be called on checkpoint. In order to restore from checkpoint, we need to know (i) which chunks have been handed out to which nodes and (ii) which data loader worker processes have yielded how many samples for each node. Mixtera implements (i) using the ChunkDistributor, which caches the query and the current state on checkpoint and is able to restore the in-memory state of the iterators for chunk generation based on this information. For (ii) the MixteraTorchDataset uses a shared memory segment to share with the main training process the status of the data loader workers. When we restore a checkpoint, we restore the state at the server, hand out the last chunks to each worker that they were working on, and then at the workers discard the previously yielded samples.
Training feedback. Dynamic mixing algorithms require a loss per property domain. Mixtera offers a simple function to forward this information to the server. Users still need to adjust the training framework. The loss implementation needs to be adjusted s.t. it is not immediately reduced but stored per domain. The per-domain losses along all training nodes need to be synchronized, e.g., via all-reduce, before passing it to Mixtera.
6. Evaluation
We evaluate Mixtera to answer the following questions:
How can we integrate dynamic mixing algorithms into Mixtera and what role do mixtures play for model accuracy?
How does Mixtera’s throughput compare to other data loaders and how well does it scale?
[…]
We explore the first question in a dynamic mixture case study on LLMs (Section 6.1) and a static mixture case study for VLMs (Section 6.2); the second question is explored in throughput benchmarks (Section 6.3). We conduct our experiments on the Clariden partition of the Alps supercomputer of the Swiss National Supercomputer Centre. Clariden consists of HPE Cray Supercomputing EX254n blades, each hosting two Quad GH200 nodes. Each node contains 4 interconnected groups of a Grace CPU with 72 cores, 128 GB DRAM, and a H100 Hopper GPU with 96 GB of HBM. The nodes are connected using 200 Gb/s HPE Slingshot interconnect. We refer to Fusco et al. (Fusco et al., 2024) for a more detailed analysis of the Alps supercomputer. The machines run Ubuntu Server 24.01 LTS with kernel 5.14.21, and we build upon the NVIDIA NGC 25.01 container with Python 3.12, a nightly build of PyTorch 2.6, NVIDIA driver 550.54.15 and CUDA 12.8. We add support for Mixtera and other data loaders on our fork of torchtitan (commit d989842)222Available at https://github.com/eth-easl/torchtitan-mixtera. (Liang et al., 2024). Torchtitan is part of the Pytorch ecosystem and straightforward to set up. We use Llama3-like models whose configurations can be found in Table 2. The 162M and 1.6B models are based on Jiang et al. (Jiang et al., 2024), while 3.6B follows Meta’s Llama 3.2 model. They do not have the same parameter count as torchtitan does not tie the weight embeddings. Our training and benchmarking data is based on The Pile (Gao et al., 2020). We heuristically split long samples with more than 3000 words, with max. 10 k samples per file.
6.1. Dynamic Mixing using ADO
We first demonstrate how to implement dynamic mixing algorithms in Mixtera, taking the ADO algorithm (Section 2.1.1) as an example. We show how dynamic mixing can improve model performance during pre-training. Dynamic mixing algorithms are an active area of research (Jiang et al., 2024; Chen et al., 2024a, 2023; Li et al., 2025; Albalak et al., 2023) and Mixtera facilitates exploration.
Model Steps Mixture HellaSwag WinoGrande ARC-E ARC-C LogiQA2 Lambada (OAI) OpenBookQA PIQA SlimP. PPL Pile PPL 162M 15 k ADO 0.28 0.49 0.43 0.20 0.23 0.29 0.19 0.61 58.19 66.88 Default 0.28 0.52 0.44 0.19 0.23 0.29 0.18 0.61 59.13 69.52 Natural 0.28 0.52 0.41 0.18 0.23 0.26 0.16 0.61 57.54 69.85 30 k ADO 0.29 0.50 0.44 0.19 0.23 0.31 0.17 0.62 52.62 61.82 Default 0.28 0.52 0.44 0.20 0.23 0.34 0.16 0.62 53.55 64.39 Natural 0.29 0.54 0.44 0.19 0.24 0.31 0.16 0.63 52.78 65.03 1B 15 k ADO 0.37 0.54 0.51 0.23 0.25 0.52 0.20 0.70 30.07 32.20 Default 0.35 0.54 0.52 0.23 0.22 0.47 0.17 0.68 29.34 33.02 Natural 0.35 0.54 0.52 0.23 0.23 0.50 0.17 0.68 31.31 34.40 30 k ADO 0.40 0.55 0.55 0.26 0.23 0.54 0.23 0.71 25.89 29.05 Default 0.38 0.56 0.56 0.24 0.22 0.56 0.21 0.70 25.10 28.65 Natural 0.38 0.56 0.55 0.25 0.22 0.57 0.21 0.69 26.55 29.83 3B 15 k ADO 0.42 0.56 0.57 0.25 0.24 0.59 0.21 0.71 25.78 23.00 Default 0.38 0.54 0.52 0.25 0.24 0.50 0.20 0.68 26.21 24.73 30 k ADO 0.45 0.59 0.60 0.28 0.24 0.59 0.24 0.73 22.22 20.52 Default 0.42 0.56 0.59 0.26 0.24 0.57 0.23 0.70 22.52 21.79
Mixtera implementation. The original ADO implementation updates the current mixture after every step and samples the next batch based on this distribution. The distribution is updated based on the per-domain loss from the previous step. This does not fully match Mixtera, which can only use a new mixture when generating a new chunk. Each chunk then may yield several batches of data with the same mixture. We still send the per-domain losses on every training step at the client to update ADO’s internal state at the server. Whenever a new chunk is generated, the current mixture from ADO is queried (Algorithm 1), and the server generates a chunk according to that mixture. This introduces some slack since we do not use a new mixture on every batch. However, due to the stochastic nature of sampling, we do not find this to be an issue.
The only change needed in training clients is the implementation of a per-domain loss. For this, the loss function (e.g., cross-entropy) needs to be called without reduction, which gives a loss per token. As Mixtera provides which token belongs to which domain, we can aggregate the losses per domain. We then perform an all-reduce operation across all training nodes to get the global per-domain losses and send this to the server.
While the original implementation of ADO is fully tied to their training framework and data setup, Mixtera is agnostic to the training framework. While developing, we switched from nanotron (HuggingFace, 2025) to torchtitan (Liang et al., 2024). No changes in Mixtera were required. Only the per-domain loss module at the training framework is required. This showcases the benefit of having a system like Mixtera that decouples the mixing from the training framework.
Training setup. We base our setup on Jiang et al. (Jiang et al., 2024). In addition to their 162M and 1.6B architectures, to test how ADO scales to the latest model architecture, we try a 3.6B model (Table 2) following Llama-3.2-3B from Meta, including its tokenizer. The 162M and 1.6B models use the EleutherAI/GPT-NeoX-20B tokenizer. All models use a sequence length of 2048. For ADO, we follow the codebase and discard the first 500 steps for fitting the scaling laws, start with fitting the scaling laws at step 1 000, and then re-fit the laws every 1 000 steps with a subsampling frequency of 10. We also follow the codebase and “use the same step size” for all domains. Instead of using the count of how often a domain has been sampled to fit the scaling laws like in the paper, we follow the ADO codebase and average the total sample counts evenly across all domains. We train all models with non-strict token-level mixtures, a learning rate of 0.001 with a linear warmup for 500 steps and linear decay until the final step, and the AdamW optimizer, for 30 000 steps. For 162M/1.6B, we use 64 GPUs with a microbatch size of 32, and for 3.6B we use 128 GPUs with a microbatch size of 16, resulting in a global batch size of 2048 and total 125 B tokens. As mixtures, we test ADO, the default weights as in the DoReMi paper (Xie et al., 2024), and on 162M/1.6B, the “natural” weights by Jiang et al. (Jiang et al., 2024) which are optimized for the NeoX tokenizer.
Evaluation metrics. We follow Jiang et al. (Jiang et al., 2024) and report both downstream task performance as well as perplexity. For downstream tasks, we report performance on HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-Easy and ARC-Challenge (Clark et al., 2018), LogiQA2 (Liu et al., 2023), Lambada OpenAI (Paperno et al., 2016), OpenBookQA (Mihaylov et al., 2018), and PIQA (Bisk et al., 2020). For perplexity, we report the average unweighted token perplexity on (i) the validation set of The Pile (Gao et al., 2020), and on (ii) SlimPajama (Soboleva et al., 2023) as a dataset we did not train on. We collect all metrics using EleutherAI’s lm-eval-harness (Gao et al., 2023), and use the unnormalized accuracy.
ADO algorithm performance overview. Table 3 shows the performance of all models and mixtures for a checkpoint after 15 k and 30 k steps. We mark in bold the best value within a model/step group. In contrast to Jiang et al. (Jiang et al., 2024), we find that on the 162M model, the default mixture is quite competitive across benchmarks. For the final checkpoint (30 k), ADO achieves the best (lowest) perplexity on both SlimPajama and The Pile, but it does not consistently beat the static mixtures as reported in the ADO paper. However, analysis of the textual outputs reveals that the quality of generated text from this small model is limited across all configurations.
For the larger 1B model, ADO is particulary strong on the intermediate checkpoint, which shows an improvement in time-to-accuracy. For the final checkpoint, on some benchmarks, the static mixtures perform on-par or marginally better than ADO, but on others, such as HellaSwag, ADO clearly beats the static mixtures.
On the 3B model, ADO beats the static default mixture on both checkpoints across all benchmarks. While Jiang et al. (Jiang et al., 2024) report that ADO performs best on the smaller scale model, we find that ADO performs better on larger models. Since ADO’s official repository is tightly coupled with the training framework, even after corresponding with the authors, we were not able to identify the root cause of this, partly also because their code is bound to training on specific cloud instances. This shows that a common codebase for data mixing decoupled from the training framework is beneficial and help developers integrate the latest algorithms into their setups (Section 3).
TODO!
Performance over time. To demonstrate how different tasks behave over the training, we show the performance of the 1B model on HellaSwag, OpenBookQA, and ARC-Easy for all training checkpoints in Figure 5. Every benchmark exhibits different behavior. For HellaSwag, we can clearly see that ADO consistently increases its margin over the static mixtures. For OpenBookQA, ADO performs well especially in the intermediate checkpoints, and the static mixtures greatly improve towards the end. For ARC-Easy, all mixtures behave similarly. This motivates future research on data mixing using Mixtera. For example, we might be able to use intermediate evaluations instead of loss to dynamically adjust the mixture.
TODO!
Mixture over time. We showcase the mixture over time for the six largest domains in Figure 6. Around step 21 k, we run out of data for the Books3 domain, and Mixtera’s best-effort algorithm re-distributes the weight proportionally to the other domains. This happens because The Pile’s samples are very imbalanced in size, i.e., the average sample in Books3 has around 538 KiB, while the average Pile-CC sample has 4 KiB. While we split long samples heuristically, there still is some waste due to token-level mixtures. Unlike Jiang et al. (Jiang et al., 2024), we do not observe that other models/tokenizers lead to different mixtures. Notably, the more parameters, the higher the weight of Books3, and the lower the weight of PubMed Central. On the 3.6B model, OpenWebText2’s weight even surpasses PubMed Central’s weight before step 5 000.
Takeaways. Dynamic mixture algorithms can improve model accuracy. ADO scales beyond the 1.6B model tested in the original paper, beating the default weights on the 3.6B model on all benchmarks. Our experiments also demonstrate that a synchronous algorithm, which uses a new mixture at each training step, adapts to Mixtera’s chunking system. We observe some discrepancies on the 162M model, which highlights the importance of having a common, open-source data mixing platform like Mixtera that facilities debugging and reproducibility.
6.2. Multimodal LLaVA Finetuning
We are not aware of prior work on the impact of data mixtures on VLMs. To showcase Mixtera’s multimodal capabilities, we evaluate the impact of static mixtures for finetuning a LLaVA-style model (Liu et al., 2024). The LLaVA framework trains an adapter between a pre-trained image encoder and a pre-trained LLM, and then fine-tunes the adapter and LLM on visual instruction-following data.
Training setup. We base our training setup on the TinyLlaVA Factory codebase by Jia et al. (Jia et al., 2024). We rely on a recipe from the Factory and use google/siglip-so400m-patch14-384 (Alabdulmohsin et al., 2023) as the vision encoder, TinyLlama/TinyLlama-1.1B-Chat-v1.0 (Zhang et al., 2024) as the LLM, and a 2-layer MLP as the adapter (Team, 2024a). We train all models on one GH200 node with 4 data parallel GPUs. For pre-training, we use a global batch size of 512 with a learning rate of 0.001, and for finetuning use a global batch size of 128 with a learning rate of 0.00002. We use a chunk size of 256, a cosine learning rate scheduler and the Adam optimizer. We follow Liu et al. (Liu et al., 2024; Team, 2023) and pre-train the adapter on a 558 k subset of the LAION-CC-SBU dataset with BLIP captions. For finetuning, we follow the TinyLLava Factory “LLaVA dataset” and use 665 k samples from six datasets (COCO (Lin et al., 2014), GQA (Hudson and Manning, 2019), OCR-VQA (Mishra et al., 2019), TextVQA (Singh et al., 2019), and VisualGenome (VG) (Krishna et al., 2017), and LLaVA’s text-only SFT annotations (Liu et al., 2024)) (Team, 2024b). We pre-train the adapter once, and then vary the proportions of the finetuning datasets. We randomly generate 256 mixtures for finetuning. Since the number of datapoints in comparison to LLM training is small, we use a best-effort mixture and ensure we go through all samples exactly once. All models see the same data, but in different order.
Benchmarks. We evaluate the models on the GQA (Hudson and Manning, 2019), SQA-IMG (Lu et al., 2022), TextVQA (Singh et al., 2019), POPE (Li et al., 2023), MME (Fu et al., 2024), and MMMU (Yue et al., 2024) benchmarks. We collect all metrics using TinyLLaVA Factory.
Results. In Table 4, we show the results reported by Jia et al. (Jia et al., 2024), the results we obtain using the inferring mixture (Section 5.2.2), and three mixtures out of the generated mixtures that perform well. While the inferring mixture does not achieve their reported results, this could be either due to different data dynamics in their training, or due to a different evaluation setup we cannot reproduce as their model weights are not public. Nevertheless, in particular on MME/MMMU/POPE, the mixtures outperform the baseline. Mix. #252 is weighted towards TextVQA (29.1 %) and OCR-VQA (19.5 %), with equal proportions of COCO (21.9%) and VG (21.9 %); GQA (4.0 %) and SFT annotations (3.6 %) contribute minimally. Mix. #107 places a large emphasis on SFT annotations (28.9 %) and VG (27.4 %), followed by COCO (25.5 %); GQA (5.3 %) and TextVQA (5.1 %) have lower representation, while OCR-VQA (7.8 %) remains a minor component. Mix. #155 prioritizes OCR-VQA (30.3 %) and TextVQA (25.7 %), and SFT annotations (22.7 %); VG (13.2 %) and COCO (6.9 %) are underrepresented, while GQA (1.2 %) is least utilized. Overall, despite seeing the same samples globally, the mixtures play a big role for model accuracy. Exploring mixtures for pre-training of VLMs and adapting dynamic mixtures to VLMs are promising future work directions.
Model SQA-IMG TxtVQA GQA MME MMMU POPE Jia et al. (Jia et al., 2024) 64.0 49.6 58.6 1256.5 28.3 86.3 Infer. Mix. 55.88 37.92 54.63 1238.76 29.4 86.6 Mix. #252 63.01 43.87 56.14 1268.05 30.4 85.7 Mix. #107 58.50 45.18 58.36 1283.62 30.1 85.54 Mix. #155 57.14 42.37 57.14 1290.06 29.3 86.63
6.3. Throughput Benchmarks
We now focus on throughput and compare Mixtera with other data loaders across various configurations. Mixtera’s goal is to avoid data stalls, i.e., training throughput should not be reduced because we are waiting for data (Mohan et al., 2021; Murray et al., 2021; Kuchnik et al., 2022; Zhao et al., 2022). Hence, we measure throughput in tokens per second in actual workloads. Note that smaller models and more data parallelism increase the pressure on the data loader, while larger models reduce the pressure. If a data loader can sustain training small models at scale on a high-end platform like the GH200, its performance is sufficient for other scenarios as well.
Besides Mixtera with chunk sizes of 512, 1024, and 2048, we benchmark the iterable HuggingFaceDataset by Torchtitan (Hf-Iter), a mapped version (Hf-Map), the Mosaic StreamingDataset (Team, 2022) (Mosaic). The difference between Hf-Iter and Hf-Map is that similar to Mixtera, Hf-Iter loads and tokenizes the data on the fly, while Hf-Map preprocesses all data, including tokenization. We evaluate throughput on the 162M model since larger models only lead to lower throughput. We always use FSDP since Torchtitan only enables bfloat16 training if sharding is used. We activate compilation, disable activation checkpointing, and use fused AdamW. We measure throughput for 30 steps, discarding the first step. We repeat all measurements three times, i.e., in total we have 3x30 steps. We use the huggingface EleutherAI/gpt-neox-20b tokenizer for all data loaders. We store all data on the Iopsstor SSD-backed Lustre DFS in the Alps cluster.
TODO!
Single-node. We train on a single GH200 node with 4 GPUs. We use 2 data parallel replicates and shards. In Figure 7 we show the throughput for the data loaders depending on the number of data workers. We test up to 16 workers, where 0 workers indicate that data is loaded in the same process as the main training loop. We only show results up to 4 workers due as throughput does not increase further. The error bars show the standard deviation of the throughput. Without data workers, Mixtera has the highest average throughput; with one worker, the other data loaders reach their peak performance.
Mixtera performs worse than the other loaders with a lower number of workers due to the random access into the files. For every sample, it needs to open the file, seek to the correct position, and load the data, instead of bulk-transferring all the data as the other data loaders can do. This leads to the higher variance in throughput indicated by the error bars, which overall leads to a lower average. However, with a higher number of workers, the variance in data access gets hidden. When using 4+ workers, Mixtera performs almost identical to the other data loaders. Additionally, increasing the chunk size also helps, and for the 0 worker case, Mixtera with a chunk size of 1024 or 2048 even has a higher average throughput.
This benchmark setup is quite extreme, as we train a very small 162M model on an extremely fast GPU. Nevertheless, just by adding some workers, which comes at no overhead beyond some CPU resources, Mixtera reaches competitive throughput despite its more complicated model.
Scaling out. We investigate how the data loaders scale for larger training jobs with higher data parallelism. We scale up to 64 GH200 nodes with a total of 256 data parallel GPUs. We find that using a maximum number of 16 replication GPUs works best. For 4, 8, and 16 GPUs, we use half the GPUs as replication, and shard across the rest. The results with 8 data workers can be found in Figure 8. All data loaders scale perfectly linear. For larger number of GPUs, the throughput variance increases a bit for all systems. We attribute this to the random assignment of nodes in the cluster by the slurm scheduler across the 3 repetitions. We do not test pipeline or tensor parallelism as (i) this is not necessary for the 162M model (and a larger model would only stress the data loaders less), (ii) the data loaders besides Mixtera do not easily support 3D parallelism, (iii) increasing data parallelism increases the load on the system more.
File formats. All previous benchmarks use uncompressed jsonl data. We test compressed jsonl (jsonl.zst), parquet, and the webdatasets format in the data loaders that support them. Notably, only Mixtera supports all of these formats. We find that the underlying file format does not impact the training throughput and therefore omit a plot. This observation holds across different numbers of data workers, where we also observe minimal performance variation among the data loaders.
TODO!
6.3.1. Query execution
Mixtera executes the query in DuckDB and generates the ChunkerIndex before we can start training. On our Pile dataset with 241 M samples, DuckDB takes 29 s, and preparing the index takes 18 s due to our C++ implementation. Optionally, we can also persist an initial state checkpoint of the query, which due to serialization of nested Python dictionaries takes 13 min. This is only necessary if checkpointing should be supported. Future checkpoints can then be stored in milliseconds. While the streaming Hf-Iter data loader can basically start streaming data immediately, the Hf-Map data loader, the default in nanotron, loads and tokenizes the data first, taking 2 h 51 min for this dataset. However, what needs to be considered is that for streaming data loaders such as Hf-Iter or Mosaic, in many scenarios, users would also need to run more offline preprocessing, e.g., to perform static filtering, or reshuffling the data if we want to mix on a different property. Mixtera avoids this offline preprocessing completely, and, besides an optional state checkpoint, starts streaming data in under a minute.
7. Conclusion and Future Work
Managing training data at scale requires data management support. We present Mixtera, a training framework-agnostic data plane for foundation model training. Using its chunking mechanism, it supports dynamic mixtures across arbitrary properties. We demonstrate Mixtera’s performance and scalability, as well as the importance of mixtures on both LLMs and VLMs. For future work, it is interesting to extend the analysis to the cloud, i.e., using a service like S3 as the underlying storage layer. Mixtera also lays the foundation to implement lineage tracking on which model was trained on which data, as it is a centralized access point. From a ML perspective, the impact of data composition and quality on model performance is still under-explored, and Mixtera provides a foundation to conduct future research on dynamic and static mixing algorithms.
Acknowledgements.
We thank Beste Güney for her contributions to Mixtera’s codebase. We thank Antoni-Joan Solergibert i Llaquet, Imanol Schlag, Steven Hand, Martin Jaggi, Antoine Bosselut, Alexander Hägele, Loubna Ben Allal, Quentin de Laroussilhe, Paul Barham, Yiding Jiang, Theodoros Rekatsinas, and Foteini Strati for helpful discussions. This work was supported as part of the Swiss AI Initiative by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID a09 on Alps. Maximilian Böther is supported by the Swiss National Science Foundation (project number 200021_204620).References
- (1)
- Akidau et al. (2015) Tyler Akidau, Robert Bradshaw, Craig Chambers, Slava Chernyak, Rafael J. Fernández-Moctezuma, Reuven Lax, Sam McVeety, Daniel Mills, Frances Perry, Eric Schmidt, and Sam Whittle. 2015. The dataflow model: a practical approach to balancing correctness, latency, and cost in massive-scale, unbounded, out-of-order data processing. Proceedings of the VLDB Endowment 8, 12 (2015). doi:10.14778/2824032.2824076
- Alabdulmohsin et al. (2023) Ibrahim M. Alabdulmohsin, Xiaohua Zhai, Alexander Kolesnikov, and Lucas Beyer. 2023. Getting ViT in Shape: Scaling Laws for Compute-Optimal Model Design. In Proceedings of the Conference on Neural Information Processing Systems (NeurIPS).
- Albalak et al. (2023) Alon Albalak, Liangming Pan, Colin Raffel, and William Yang Wang. 2023. Efficient Online Data Mixing For Language Model Pre-Training. arXiv Preprint (2023). doi:10.48550/arXiv.2312.02406
- Allal et al. (2025) Loubna Ben Allal, Anton Lozhkov, Elie Bakouch, Gabriel Martín Blázquez, Guilherme Penedo, Lewis Tunstall, Andrés Marafioti, Hynek Kydlíček, Agustín Piqueres Lajarín, Vaibhav Srivastav, Joshua Lochner, Caleb Fahlgren, Xuan-Son Nguyen, Clémentine Fourrier, Ben Burtenshaw, Hugo Larcher, Haojun Zhao, Cyril Zakka, Mathieu Morlon, Colin Raffel, Leandro von Werra, and Thomas Wolf. 2025. SmolLM2: When Smol Goes Big – Data-Centric Training of a Small Language Model. arXiv preprint (2025). doi:10.48550/arXiv.2502.02737
- Armbrust et al. (2021) Michael Armbrust, Ali Ghodsi, Reynold Xin, and Matei Zaharia. 2021. Lakehouse: A New Generation of Open Platforms that Unify Data Warehousing and Advanced Analytics. In Proceedings of the Conference on Innovative Data Systems Research (CIDR).
- Ben-Nun and Hoefler (2019) Tal Ben-Nun and Torsten Hoefler. 2019. Demystifying Parallel and Distributed Deep Learning: An In-depth Concurrency Analysis. Comput. Surveys 52, 4 (2019), 1–43. doi:10.1145/3320060
- Bisk et al. (2020) Yonatan Bisk, Rowan Zellers, Ronan Le Bras, Jianfeng Gao, and Yejin Choi. 2020. PIQA: Reasoning about Physical Commonsense in Natural Language. Proceedings of the Conference on Artificial Intelligence (AAAI). doi:10.1609/aaai.v34i05.6239
- Bommasani et al. (2022) Rishi Bommasani, Drew A. Hudson, Ehsan Adeli, Russ Altman, Simran Arora, Sydney von Arx, Michael S. Bernstein, Jeannette Bohg, Antoine Bosselut, Emma Brunskill, Erik Brynjolfsson, Shyamal Buch, Dallas Card, Rodrigo Castellon, Niladri Chatterji, Annie Chen, Kathleen Creel, Jared Quincy Davis, Dora Demszky, Chris Donahue, Moussa Doumbouya, Esin Durmus, Stefano Ermon, John Etchemendy, Kawin Ethayarajh, Li Fei-Fei, Chelsea Finn, Trevor Gale, Lauren Gillespie, Karan Goel, Noah Goodman, Shelby Grossman, Neel Guha, Tatsunori Hashimoto, Peter Henderson, John Hewitt, Daniel E. Ho, Jenny Hong, Kyle Hsu, Jing Huang, Thomas Icard, Saahil Jain, Dan Jurafsky, Pratyusha Kalluri, Siddharth Karamcheti, Geoff Keeling, Fereshte Khani, Omar Khattab, Pang Wei Koh, Mark Krass, Ranjay Krishna, Rohith Kuditipudi, Ananya Kumar, Faisal Ladhak, Mina Lee, Tony Lee, Jure Leskovec, Isabelle Levent, Xiang Lisa Li, Xuechen Li, Tengyu Ma, Ali Malik, Christopher D. Manning, Suvir Mirchandani, Eric Mitchell, Zanele Munyikwa, Suraj Nair, Avanika Narayan, Deepak Narayanan, Ben Newman, Allen Nie, Juan Carlos Niebles, Hamed Nilforoshan, Julian Nyarko, Giray Ogut, Laurel Orr, Isabel Papadimitriou, Joon Sung Park, Chris Piech, Eva Portelance, Christopher Potts, Aditi Raghunathan, Rob Reich, Hongyu Ren, Frieda Rong, Yusuf Roohani, Camilo Ruiz, Jack Ryan, Christopher Ré, Dorsa Sadigh, Shiori Sagawa, Keshav Santhanam, Andy Shih, Krishnan Srinivasan, Alex Tamkin, Rohan Taori, Armin W. Thomas, Florian Tramèr, Rose E. Wang, William Wang, Bohan Wu, Jiajun Wu, Yuhuai Wu, Sang Michael Xie, Michihiro Yasunaga, Jiaxuan You, Matei Zaharia, Michael Zhang, Tianyi Zhang, Xikun Zhang, Yuhui Zhang, Lucia Zheng, Kaitlyn Zhou, and Percy Liang. 2022. On the Opportunities and Risks of Foundation Models. In arXiv preprint. doi:10.48550/arXiv.2108.07258
- Brown et al. (2020) Tom B. Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel M. Ziegler, Jeffrey Wu, Clemens Winter, Christopher Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020. Language Models are Few-Shot Learners. In Proceedings of Advances in Neural Information Processing Systems (NeurIPS).
- Chen et al. (2024b) Daoyuan Chen, Yilun Huang, Zhijian Ma, Hesen Chen, Xuchen Pan, Ce Ge, Dawei Gao, Yuexiang Xie, Zhaoyang Liu, Jinyang Gao, Yaliang Li, Bolin Ding, and Jingren Zhou. 2024b. Data-Juicer: A One-Stop Data Processing System for Large Language Models. In Proceedings of the International Conference on Management of Data (SIGMOD). doi:10.1145/3626246.3653385
- Chen et al. (2024a) Mayee F. Chen, Michael Y. Hu, Nicholas Lourie, Kyunghyun Cho, and Christopher Ré. 2024a. Aioli: A Unified Optimization Framework for Language Model Data Mixing. In Proceedings of the International Conference on Learning Representations (ICLR). doi:10.48550/arXiv.2411.05735
- Chen et al. (2023) Mayee F. Chen, Nicholas Roberts, Kush Bhatia, Jue Wang, Ce Zhang, Frederic Sala, and Christopher Ré. 2023. Skill-it! A Data-Driven Skills Framework for Understanding and Training Language Models. In Proceedings of Advances in Neural Information Processing Systems (NeurIPS). doi:10.48550/arXiv.2307.14430
- Clark et al. (2018) Peter Clark, Isaac Cowhey, Oren Etzioni, Tushar Khot, Ashish Sabharwal, Carissa Schoenick, and Oyvind Tafjord. 2018. Think you have Solved Question Answering? Try ARC, the AI2 Reasoning Challenge. arXiv preprint (2018). doi:10.48550/arXiv.1803.05457
- Competition and Markets Authority (2013) Competition and Markets Authority. 2013. AI Foundation Models: Initial Report. Technical Report. UK Government Agency.
- Cooper et al. (2022) A. Feder Cooper, Jonathan Frankle, and Christopher De Sa. 2022. Non-Determinism and the Lawlessness of Machine Learning Code. In Proceedings of the Symposium on Computer Science and Law (CSLAW). doi:10.1145/3511265.3550446
- ETH AI Center & EPFL AI Center (2025) ETH AI Center & EPFL AI Center. 2025. Swiss AI Initiative. https://www.swiss-ai.org/.
- Fu et al. (2024) Chaoyou Fu, Peixian Chen, Yunhang Shen, Yulei Qin, Mengdan Zhang, Xu Lin, Jinrui Yang, Xiawu Zheng, Ke Li, Xing Sun, Yunsheng Wu, and Rongrong Ji. 2024. MME: A Comprehensive Evaluation Benchmark for Multimodal Large Language Models. arXiv preprint (2024). doi:10.48550/ARXIV.2306.13394
- Fusco et al. (2024) Luigi Fusco, Mikhail Khalilov, Marcin Chrapek, Giridhar Chukkapalli, Thomas C. Schulthess, and Torsten Hoefler. 2024. Understanding Data Movement in Tightly Coupled Heterogeneous Systems: A Case Study with the Grace Hopper Superchip. arXiv preprint (2024). doi:10.48550/ARXIV.2408.11556
- Gao et al. (2020) Leo Gao, Stella Biderman, Sid Black, Laurence Golding, Travis Hoppe, Charles Foster, Jason Phang, Horace He, Anish Thite, Noa Nabeshima, Shawn Presser, and Connor Leahy. 2020. The Pile: An 800GB Dataset of Diverse Text for Language Modeling. arXiv preprint (2020). doi:10.48550/arXiv.2101.00027
- Gao et al. (2023) Leo Gao, Jonathan Tow, Baber Abbasi, Stella Biderman, Sid Black, Anthony DiPofi, Charles Foster, Laurence Golding, Jeffrey Hsu, Alain Le Noac’h, Haonan Li, Kyle McDonell, Niklas Muennighoff, Chris Ociepa, Jason Phang, Laria Reynolds, Hailey Schoelkopf, Aviya Skowron, Lintang Sutawika, Eric Tang, Anish Thite, Ben Wang, Kevin Wang, and Andy Zou. 2023. A framework for few-shot language model evaluation. doi:10.5281/zenodo.10256836
- Hoefler et al. (2024) Torsten Hoefler, Tommaso Bonoto, Daniele De Sensi, Salvatore Di Girolamo, Shigang Li, Marco Heddes, Deepak Goel, Miguel Castro, and Steve Scott. 2024. HammingMesh: A Network Topology for Large-Scale Deep Learning. Commun. ACM 67, 12 (2024), 97–105. doi:10.1145/3623490
- Huang et al. (2019) Yanping Huang, Youlong Cheng, Ankur Bapna, Orhan Firat, Dehao Chen, Mia Xu Chen, HyoukJoong Lee, Jiquan Ngiam, Quoc V. Le, Yonghui Wu, and Zhifeng Chen. 2019. GPipe: Efficient Training of Giant Neural Networks using Pipeline Parallelism. In Proceedings of Advances in Neural Information Processing Systems (NeurIPS).
- Hudson and Manning (2019) Drew A. Hudson and Christopher D. Manning. 2019. GQA: A New Dataset for Real-World Visual Reasoning and Compositional Question Answering. In Proceedings of the Conference on Computer Vision and Pattern Recognition (CVPR). doi:10.1109/cvpr.2019.00686
- HuggingFace (2025) HuggingFace. 2025. Nanotron: Pretraining models made easy. https://github.com/huggingface/nanotron
- Jia et al. (2024) Junlong Jia, Ying Hu, Xi Weng, Yiming Shi, Miao Li, Xingjian Zhang, Baichuan Zhou, Ziyu Liu, Jie Luo, Lei Huang, and Ji Wu. 2024. TinyLLaVA Factory: A Modularized Codebase for Small-scale Large Multimodal Models. arXiv preprint (2024). doi:10.48550/arXiv.2405.11788
- Jiang et al. (2024) Yiding Jiang, Allan Zhou, Zhili Feng, Sadhika Malladi, and J. Zico Kolter. 2024. Adaptive Data Optimization: Dynamic Sample Selection with Scaling Laws. In Proceedings of the International Conference on Learning Representations (ICLR). doi:10.48550/arXiv.2410.11820
- Karamcheti et al. (2021) Siddharth Karamcheti, Laurel Orr, Jason Bolton, Tianyi Zhang, Karan Goel, Avanika Narayan, Rishi Bommasani, Deepak Narayanan, Tatsunori Hashimoto, Dan Jurafsky, Christopher D. Manning, Christopher Potts, Christopher Ré, and Percy Liang. 2021. Mistral - A Journey towards Reproducible Language Model Training.
- Krishna et al. (2017) Ranjay Krishna, Yuke Zhu, Oliver Groth, Justin Johnson, Kenji Hata, Joshua Kravitz, Stephanie Chen, Yannis Kalantidis, Li-Jia Li, David A. Shamma, Michael S. Bernstein, and Li Fei-Fei. 2017. Visual Genome: Connecting Language and Vision Using Crowdsourced Dense Image Annotations. International Journal of Computer Vision 123, 1 (2017). doi:10.1007/s11263-016-0981-7
- Kuchnik et al. (2022) Michael Kuchnik, Ana Klimovic, Jiri Simsa, Virginia Smith, and George Amvrosiadis. 2022. Plumber: Diagnosing and Removing Performance Bottlenecks in Machine Learning Data Pipelines. In Proceedings of the Conference on Machine Learning and Systems (MLSys).
- Li et al. (2023) Yifan Li, Yifan Du, Kun Zhou, Jinpeng Wang, Xin Zhao, and Ji-Rong Wen. 2023. Evaluating Object Hallucination in Large Vision-Language Models. In Proceedings of the Conference on Empirical Methods in Natural Language Processing (EMNLP). doi:10.18653/v1/2023.emnlp-main.20
- Li et al. (2025) Zeman Li, Yuan Deng, Peilin Zhong, Meisam Razaviyayn, and Vahab Mirrokni. 2025. PiKE: Adaptive Data Mixing for Multi-Task Learning Under Low Gradient Conflicts. arXiv preprint (2025). doi:10.48550/arXiv.2502.06244
- Liang et al. (2024) Wanchao Liang, Tianyu Liu, Less Wright, Will Constable, Andrew Gu, Chien-Chin Huang, Iris Zhang, Wei Feng, Howard Huang, Junjie Wang, Sanket Purandare, Gokul Nadathur, and Stratos Idreos. 2024. TorchTitan: One-stop PyTorch native solution for production ready LLM pre-training. arXiv preprint (2024). doi:10.48550/arXiv.2410.06511
- Lin et al. (2014) Tsung-Yi Lin, Michael Maire, Serge J. Belongie, James Hays, Pietro Perona, Deva Ramanan, Piotr Dollár, and C. Lawrence Zitnick. 2014. Microsoft COCO: Common Objects in Context. In Proceedings of the European Conference on Computer Vision (ECCV). doi:10.1007/978-3-319-10602-1_48
- Liu et al. (2024) Haotian Liu, Chunyuan Li, Yuheng Li, and Yong Jae Lee. 2024. Improved Baselines with Visual Instruction Tuning. In Proceedings of the Conference on Computer Vision and Pattern Recognition (CVPR). doi:10.1109/cvpr52733.2024.02484
- Liu et al. (2023) Hanmeng Liu, Jian Liu, Leyang Cui, Zhiyang Teng, Nan Duan, Ming Zhou, and Yue Zhang. 2023. LogiQA 2.0—An Improved Dataset for Logical Reasoning in Natural Language Understanding. IEEE/ACM Transactions on Audio, Speech, and Language Processing 31 (2023). doi:10.1109/taslp.2023.3293046
- Longpre et al. (2024) Shayne Longpre, Gregory Yauney, Emily Reif, Katherine Lee, Adam Roberts, Barret Zoph, Denny Zhou, Jason Wei, Kevin Robinson, David Mimno, and Daphne Ippolito. 2024. A Pretrainer’s Guide to Training Data: Measuring the Effects of Data Age, Domain Coverage, Quality, & Toxicity. In Proceedings of the Conference of the North American Chapter of the Association for Computational Linguistics (NAACL). doi:10.18653/v1/2024.naacl-long.179
- Lu et al. (2022) Pan Lu, Swaroop Mishra, Tony Xia, Liang Qiu, Kai-Wei Chang, Song-Chun Zhu, Oyvind Tafjord, Peter Clark, and Ashwin Kalyan. 2022. Learn to Explain: Multimodal Reasoning via Thought Chains for Science Question Answering. In Proceedings of the Conference on Neural Information Processing Systems (NeurIPS).
- Meta (2024a) Meta. 2024a. The Llama 3 Herd of Models. arXiv preprint (2024). doi:10.48550/arXiv.2407.21783
- Meta (2024b) Meta. 2024b. Llama 3.3 Model Card. https://github.com/meta-llama/llama-models/blob/main/models/llama3_3/MODEL_CARD.md. Accessed: 2024-12-18.
- Mihaylov et al. (2018) Todor Mihaylov, Peter Clark, Tushar Khot, and Ashish Sabharwal. 2018. Can a Suit of Armor Conduct Electricity? A New Dataset for Open Book Question Answering. In Proceedings of the Conference on Empirical Methods in Natural Language Processing (EMNLP). doi:10.18653/v1/d18-1260
- Mishra et al. (2019) Anand Mishra, Shashank Shekhar, Ajeet Kumar Singh, and Anirban Chakraborty. 2019. OCR-VQA: Visual Question Answering by Reading Text in Images. In Proceedings of the International Conference on Document Analysis and Recognition (ICDAR). doi:10.1109/icdar.2019.00156
- Mohan et al. (2021) Jayashree Mohan, Amar Phanishayee, Ashish Raniwala, and Vijay Chidambaram. 2021. Analyzing and mitigating data stalls in DNN training. Proceedings of the VLDB Endowment 14, 5 (2021). doi:10.14778/3446095.3446100
- Murray et al. (2021) Derek G. Murray, Jiří Šimša, Ana Klimovic, and Ihor Indyk. 2021. tf.data: a machine learning data processing framework. Proceedings of the VLDB Endowment 14, 12 (2021). doi:10.14778/3476311.3476374
- Narayanan et al. (2019) Deepak Narayanan, Aaron Harlap, Amar Phanishayee, Vivek Seshadri, Nikhil R. Devanur, Gregory R. Ganger, Phillip B. Gibbons, and Matei Zaharia. 2019. PipeDream: generalized pipeline parallelism for DNN training. In Proceedings of the Symposium on Operating Systems Principles (SOSP). doi:10.1145/3341301.3359646
- Narayanan et al. (2021) Deepak Narayanan, Mohammad Shoeybi, Jared Casper, Patrick LeGresley, Mostofa Patwary, Vijay Korthikanti, Dmitri Vainbrand, Prethvi Kashinkunti, Julie Bernauer, Bryan Catanzaro, Amar Phanishayee, and Matei Zaharia. 2021. Efficient Large-Scale Language Model Training on GPU Clusters Using Megatron-LM. In Proceedings of the International Conference for High Performance Computing, Networking, Storage and Analysis (SC). doi:10.1145/3458817.3476209
- OpenAI (2024) OpenAI. 2024. GPT-4 Technical Report. In arXiv preprint. doi:10.48550/arXiv.2303.08774
- Paperno et al. (2016) Denis Paperno, Germán Kruszewski, Angeliki Lazaridou, Ngoc Quan Pham, Raffaella Bernardi, Sandro Pezzelle, Marco Baroni, Gemma Boleda, and Raquel Fernandez. 2016. The LAMBADA dataset: Word prediction requiring a broad discourse context. In Proceedings of the Annual Meeting of the Association for Computational Linguistics (ACL). doi:10.18653/v1/p16-1144
- Penedo et al. (2024a) Guilherme Penedo, Hynek Kydlíček, Loubna Ben allal, Anton Lozhkov, Margaret Mitchell, Colin Raffel, Leandro Von Werra, and Thomas Wolf. 2024a. The FineWeb Datasets: Decanting the Web for the Finest Text Data at Scale. arXiv preprint (2024). doi:10.48550/arXiv.2406.17557
- Penedo et al. (2024b) Guilherme Penedo, Hynek Kydlíček, Alessandro Cappelli, Mario Sasko, and Thomas Wolf. 2024b. DataTrove: large scale data processing. https://github.com/huggingface/datatrove
- Qian et al. (2021) Shangshu Qian, Hung Viet Pham, Thibaud Lutellier, Zeou Hu, Jungwon Kim, Lin Tan, Yaoliang Yu, Jiahao Chen, and Sameena Shah. 2021. Are My Deep Learning Systems Fair? An Empirical Study of Fixed-Seed Training. In Proceedings of the Conference on Neural Information Processing Systems (NeurIPS).
- Raasveldt and Mühleisen (2019) Mark Raasveldt and Hannes Mühleisen. 2019. DuckDB: An Embeddable Analytical Database. In Proceedings of the International Conference on Management of Data (SIGMOD). doi:10.1145/3299869.3320212
- Radford et al. (2019) Alec Radford, Jeff Wu, Rewon Child, David Luan, Dario Amodei, and Ilya Sutskever. 2019. Language Models are Unsupervised Multitask Learners. In Self-hosted preprint.
- Richburg and Carpuat (2024) Aquia Richburg and Marine Carpuat. 2024. How Multilingual Are Large Language Models Fine-Tuned for Translation? arXiv preprint (2024). doi:10.48550/arXiv.2405.20512
- Sakaguchi et al. (2021) Keisuke Sakaguchi, Ronan Le Bras, Chandra Bhagavatula, and Yejin Choi. 2021. WinoGrande: An Adversarial Winograd Schema Challenge at Scale. Commun. ACM 64, 9 (2021). doi:10.1145/3474381
- Shazeer et al. (2018) Noam Shazeer, Youlong Cheng, Niki Parmar, Dustin Tran, Ashish Vaswani, Penporn Koanantakool, Peter Hawkins, HyoukJoong Lee, Mingsheng Hong, Cliff Young, Ryan Sepassi, and Blake A. Hechtman. 2018. Mesh-TensorFlow: Deep Learning for Supercomputers. In Proceedings of Advances in Neural Information Processing Systems (NeurIPS).
- Shen et al. (2024) Zhiqiang Shen, Tianhua Tao, Liqun Ma, Willie Neiswanger, Zhengzhong Liu, Hongyi Wang, Bowen Tan, Joel Hestness, Natalia Vassilieva, Daria Soboleva, and Eric Xing. 2024. SlimPajama-DC: Understanding Data Combinations for LLM Training. arXiv preprint (2024). 10.48550/arXiv.2309.10818
- Shoeybi et al. (2019) Mohammad Shoeybi, Mostofa Patwary, Raul Puri, Patrick LeGresley, Jared Casper, and Bryan Catanzaro. 2019. Megatron-LM: Training Multi-Billion Parameter Language Models Using Model Parallelism. arXiv preprint (2019). doi:10.48550/arXiv.1909.08053
- Singh et al. (2019) Amanpreet Singh, Vivek Natarajan, Meet Shah, Yu Jiang, Xinlei Chen, Dhruv Batra, Devi Parikh, and Marcus Rohrbach. 2019. Towards VQA Models That Can Read. In Proceedings of the Conference on Computer Vision and Pattern Recognition (CVPR). doi:10.1109/cvpr.2019.00851
- Soboleva et al. (2023) Daria Soboleva, Faisal Al-Khateeb, Robert Myers, Jacob R Steeves, Joel Hestness, and Nolan Dey. 2023. SlimPajama: A 627B token cleaned and deduplicated version of RedPajama. https://huggingface.co/datasets/cerebras/SlimPajama-627B
- Soldaini et al. (2024) Luca Soldaini, Rodney Kinney, Akshita Bhagia, Dustin Schwenk, David Atkinson, Russell Authur, Ben Bogin, Khyathi Chandu, Jennifer Dumas, Yanai Elazar, Valentin Hofmann, Ananya Harsh Jha, Sachin Kumar, Li Lucy, Xinxi Lyu, Nathan Lambert, Ian Magnusson, Jacob Morrison, Niklas Muennighoff, Aakanksha Naik, Crystal Nam, Matthew E. Peters, Abhilasha Ravichander, Kyle Richardson, Zejiang Shen, Emma Strubell, Nishant Subramani, Oyvind Tafjord, Pete Walsh, Luke Zettlemoyer, Noah A. Smith, Hannaneh Hajishirzi, Iz Beltagy, Dirk Groeneveld, Jesse Dodge, and Kyle Lo. 2024. Dolma: An Open Corpus of Three Trillion Tokens for Language Model Pretraining Research. arXiv preprint (2024). doi:10.48550/arXiv.2402.00159
- Team (2023) The LlaVA Team. 2023. LLaVA Data Documentation. https://github.com/haotian-liu/LLaVA/blob/main/docs/Data.md
- Team (2024a) The LlaVA Team. 2024a. TinyLLaVA: Model Zoo. https://github.com/TinyLLaVA/TinyLLaVA_Factory?tab=readme-ov-file#model-zoo
- Team (2022) The Mosaic ML Team. 2022. streaming: Fast, accurate streaming of training data from cloud storage. https://github.com/mosaicml/streaming/
- Team (2025) The TensorFlow Team. 2025. Tensorflow: Determinism. https://www.tensorflow.org/api_docs/python/tf/config/experimental/enable_op_determinism
- Team (2024b) The TinyLLaVA Factory Team. 2024b. TinyLLaVA Factory: Prepare Datasets. https://tinyllava-factory.readthedocs.io/en/latest/Prepare%20Datasets.html
- Vaswani et al. (2017) Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan N. Gomez, Lukasz Kaiser, and Illia Polosukhin. 2017. Attention is All you Need. In Proceedings of Advances in Neural Information Processing Systems (NeurIPS).
- Wang et al. (2023) Zige Wang, Wanjun Zhong, Yufei Wang, Qi Zhu, Fei Mi, Baojun Wang, Lifeng Shang, Xin Jiang, and Qun Liu. 2023. Data Management For Large Language Models: A Survey. arXiv preprint (2023). doi:10.48550/ARXIV.2312.01700
- Weber et al. (2024) Maurice Weber, Daniel Y Fu, Quentin Gregory Anthony, Yonatan Oren, Shane Adams, Anton Alexandrov, Xiaozhong Lyu, Huu Nguyen, Xiaozhe Yao, Virginia Adams, Ben Athiwaratkun, Rahul Chalamala, Kezhen Chen, Max Ryabinin, Tri Dao, Percy Liang, Christopher Re, Irina Rish, and Ce Zhang. 2024. RedPajama: an Open Dataset for Training Large Language Models. In Proceedings of Advances in Neural Information Processing Systems (NeurIPS).
- Xie et al. (2024) Sang Michael Xie, Hieu Pham, Xuanyi Dong, Nan Du, Hanxiao Liu, Yifeng Lu, Percy S Liang, Quoc V Le, Tengyu Ma, and Adams Wei Yu. 2024. Doremi: Optimizing data mixtures speeds up language model pretraining. Advances in Neural Information Processing Systems 36 (2024).
- Xu et al. (2024b) Canwen Xu, Corby Rosset, Ethan C. Chau, Luciano Del Corro, Shweti Mahajan, Julian McAuley, Jennifer Neville, Ahmed Hassan Awadallah, and Nikhil Rao. 2024b. Automatic Pair Construction for Contrastive Post-training. arXiv preprint (2024). 10.48550/arXiv.2310.02263
- Xu et al. (2024a) Haoran Xu, Young Jin Kim, Amr Sharaf, and Hany Hassan Awadalla. 2024a. A Paradigm Shift in Machine Translation: Boosting Translation Performance of Large Language Models. In Proceedings of the ML Evaluation Standards Workshop at ICLR. doi:10.48550/arXiv.2309.1167
- Ye et al. (2024) Jiasheng Ye, Peiju Liu, Tianxiang Sun, Yunhua Zhou, Jun Zhan, and Xipeng Qiu. 2024. Data Mixing Laws: Optimizing Data Mixtures by Predicting Language Modeling Performance. arXiv preprint (2024). doi:10.48550/arXiv.2403.16952
- Yue et al. (2024) Xiang Yue, Yuansheng Ni, Tianyu Zheng, Kai Zhang, Ruoqi Liu, Ge Zhang, Samuel Stevens, Dongfu Jiang, Weiming Ren, Yuxuan Sun, Cong Wei, Botao Yu, Ruibin Yuan, Renliang Sun, Ming Yin, Boyuan Zheng, Zhenzhu Yang, Yibo Liu, Wenhao Huang, Huan Sun, Yu Su, and Wenhu Chen. 2024. MMMU: A Massive Multi-Discipline Multimodal Understanding and Reasoning Benchmark for Expert AGI. In Proceedings of the Conference on Computer Vision and Pattern Recognition (CVPR). doi:10.1109/cvpr52733.2024.00913
- Zaharia et al. (2016) Matei Zaharia, Reynold S. Xin, Patrick Wendell, Tathagata Das, Michael Armbrust, Ankur Dave, Xiangrui Meng, Josh Rosen, Shivaram Venkataraman, Michael J. Franklin, Ali Ghodsi, Joseph Gonzalez, Scott Shenker, and Ion Stoica. 2016. Apache Spark: a unified engine for big data processing. Commun. ACM 59, 11 (2016), 56–65. doi:10.1145/2934664
- Zellers et al. (2019) Rowan Zellers, Ari Holtzman, Yonatan Bisk, Ali Farhadi, and Yejin Choi. 2019. HellaSwag: Can a Machine Really Finish Your Sentence?. In Proceedings of the Annual Meeting of the Association for Computational Linguistics (ACL). doi:10.18653/v1/p19-1472
- Zhang et al. (2024) Peiyuan Zhang, Guangtao Zeng, Tianduo Wang, and Wei Lu. 2024. TinyLlama: An Open-Source Small Language Model. arXiv preprint (2024). doi:10.48550/arXiv.2401.02385
- Zhao et al. (2022) Mark Zhao, Niket Agarwal, Aarti Basant, Buğra Gedik, Satadru Pan, Mustafa Ozdal, Rakesh Komuravelli, Jerry Pan, Tianshu Bao, Haowei Lu, Sundaram Narayanan, Jack Langman, Kevin Wilfong, Harsha Rastogi, Carole-Jean Wu, Christos Kozyrakis, and Parik Pol. 2022. Understanding Data Storage and Ingestion for Large-Scale Deep Recommendation Model Training. In Proceedings of the Annual International Symposium on Computer Architecture (ISCA). doi:10.1145/3470496.3533044
- Zhao et al. (2023) Yanli Zhao, Andrew Gu, Rohan Varma, Liang Luo, Chien-Chin Huang, Min Xu, Less Wright, Hamid Shojanazeri, Myle Ott, Sam Shleifer, Alban Desmaison, Can Balioglu, Pritam Damania, Bernard Nguyen, Geeta Chauhan, Yuchen Hao, Ajit Mathews, and Shen Li. 2023. PyTorch FSDP: Experiences on Scaling Fully Sharded Data Parallel. Proceedings of the VLDB Endowment 16, 12 (2023). doi:10.14778/3611540.3611569
- Zhuang et al. (2022) Donglin Zhuang, Xingyao Zhang, Shuaiwen Song, and Sara Hooker. 2022. Randomness in Neural Network Training: Characterizing the Impact of Tooling. In Proceedings of the Conference on Machine Learning and Systems (MLSys).

### v3

@@
 We experimentally evaluate Mixtera and show that our implementation does not bottleneck training and scales to 256 GH200 superchips.
-We demonstrate how Mixtera supports recent advancements in mixing strategies by implementing the proposed Adaptive Data Optimization (ADO) algorithm in the system and evaluating its performance impact.
-We also explore the role of mixtures for vision-language models.
-PVLDB Reference Format:
-PVLDB, 1(1): 1 - 1, 1.
-doi:1
-††This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
-Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
-For any use beyond those covered by this license, obtain permission by emailing info@vldb.org.
-Copyright is held by the owner/author(s).
-Publication rights licensed to the VLDB Endowment.
-Proceedings of the VLDB Endowment, Vol. 1, No. 1 ISSN 2150-8097.
-doi:1
-PVLDB Artifact Availability:
-The source code, data, and/or other artifacts have been made available at https://github.com/eth-easl/mixtera.
+We demonstrate how Mixtera supports recent advancements in mixing strategies by implementing the Adaptive Data Optimization (ADO) algorithm in the system and evaluating its performance impact.
+We also show how Mixtera enables exploring the role of mixtures for vision-language models, which is a growing area of research.
 1. Introduction
-Large language and vision models (often called foundation models) have become omnipresent in our daily life.
+Large language and vision models (LLMs/VLMs, often called foundation models) have become omnipresent in our daily lives.
 They show enormous capabilities in a diverse set of tasks (Brown et al., 2020; Bommasani et al., 2022; OpenAI, 2024; Meta, 2024a; Radford et al., 2019), such as assistance with writing and coding, video understanding, and even agentic interaction with the world.
-The training of such language and vision models (LLMs/VLMs) presents new challenges for managing training data due to the ever-growing sizes of models and datasets.
-To achieve high accuracy, state-of-the-art models train over trillions of tokens from aggregated data collections such as RedPajama (Weber et al., 2024), Dolma (Soldaini et al., 2024), or FineWeb (Penedo et al., 2024a).
+The training of such language and vision models presents new challenges for managing training data due to the ever-growing sizes of models and datasets.
+To achieve high accuracy, state-of-the-art models train over trillions of tokens.
 For example, Meta’s Llama 3.3 70B model is trained on a corpus of 15 trillion tokens (Meta, 2024a, b).
-These collections are built based on data from various sources, such as Wikipedia or Common Crawl dumps.
-Training data is typically stored on distributed filesystems in GPU clusters or data lakes in the cloud, and managed manually and ad hoc by ML engineers (Figure 2a).
-Selecting the right proportions out of this data with particular characteristics (e.g., language, topic, source) for training is critical for model performance (Chen et al., 2023; Ye et al., 2024; Xie et al., 2024).
-Individual users write ad hoc scripts to process the training data, filter relevant subsets with the properties of interest, often pre-tokenize it, and then mix it for their use case, e.g., they might want to train on 50 % data from Wikipedia and 50 % from movie subtitles.
-This reflects our experience working with ML researchers as part of the Swiss AI initiative which aims to develop open-source LLMs (ETH AI Center & EPFL AI Center, 2025).
-We confirmed through discussions with industry teams that such setups are common.
-This can get complex as training data may need to be mixed based on a variety of characteristics.
-For example, in addition to satisfying source data proportions (Wikipedia vs.
-movie subtitles in the previous example), we may also want the training data to be 80 % in English and 20 % in German.
-Recent works show that the data mixture should also be adjusted during training.
-The SmolLM2 model is trained with four stages of data mixtures (Allal et al., 2025).
-Algorithms such as Adaptive Data Optimization (ADO) (Jiang et al., 2024), Aioli (Chen et al., 2024a), PiKE (Li et al., 2025), and Skill-It (Chen et al., 2023) even propose to adjust the mixture dynamically based on the model behavior (e.g., loss per domain) during training.
-Figure 1 shows that using ADO can increase the performance of LLMs over a static mixture across different scales on the downstream HellaSwag benchmark (Zellers et al., 2019).
-Today, there is no open-source system that automatically manages vast amounts of training data based on fine-grained characteristics.
-Implementing data filtering and subsequent mixing requires users to manually keep track of metadata.
-At least in the open-source world, this is typically done as part of the directory structure of the filesystem, e.g., there is one subdirectory per source dataset, and then we sample from each directory (Figure 2a).
+These tokens typically come from aggregated data collections such as RedPajama (Weber et al., 2024), Dolma (Soldaini et al., 2024), or FineWeb (Penedo et al., 2024a), which include data from various sources, such as Wikipedia or Common Crawl dumps.
+The composition of training data is critical to model quality (Chen et al., 2024b).
+Hence, selecting the right proportions of data with particular characteristics (e.g., language, topic, source) has become an active area of research to improve model performance without increasing training compute budget (Chen et al., 2023; Ye et al., 2024; Xie et al., 2024).
+For example, Hugging Face’s SmolLM2 model is trained with four stages of data mixtures that combine web text with specialized math, code, and instruction-following data in varying proportions (Allal et al., 2025).
+Algorithms such as Adaptive Data Optimization (ADO) (Jiang et al., 2024), Aioli (Chen et al., 2024a), PiKE (Li et al., 2025), and Skill-It (Chen et al., 2023) even propose adjusting the data mixture dynamically based on the model behavior (e.g., loss per domain) during training.
+Figure˜1 shows that ADO increases the accuracy of 1.6B and 3.6B Llama-models (c.f.
+Table˜2) compared to using a static mixture on the downstream HellaSwag benchmark (Zellers et al., 2019).
+However, the process of composing training data mixtures today is manual, ad hoc, and error-prone (Figure˜2a).
+Training data is typically stored on distributed filesystems in GPU clusters or data lakes in the cloud.
+ML engineers and researchers write ad hoc scripts to process the training data, filter relevant subsets with the properties of interest, often pre-tokenize it, and then mix it for their use case.
+For example, a model developer may want to train on 50 % data from Wikipedia and 50 % from movie subtitles.
+This can quickly become more complex as training data may need to be mixed based on multiple characteristics.
+For example, in addition to source data proportions (Wikipedia vs.
+movie subtitles), the developer may also want the training data to be 80 % in English and 20 % in German.
+When using mixing algorithms such as ADO, the data mixture may also need to be adjusted on the fly during training.
+Model developers currently lack an open-source solution to efficiently manage and declaratively query vast amounts of training data based on the characteristics of individual data elements.
+This reflects our experience working with ML researchers as part of a large-scale initiative developing open-source LLMs and our conversations with industry teams.
+Implementing data filtering and subsequent mixing requires today’s model developers to manually keep track of metadata.
+At least in the open-source world, developers typically implement this as part of the directory structure of the filesystem, e.g., developers create one subdirectory per source dataset and then sample from each directory (Figure˜2a).
 This approach is limited because each data sample has multiple properties that can be used to determine whether it should be used for training.
 Filesystems fundamentally do not offer the right interface for managing training data and mixing, as they do not provide declarative query interfaces or a native way to track which model was trained on what data.
-Running a data processing job to fully materialize the mixed training set for each training run has lots of overhead and leads to data duplication.
-A streaming-oriented approach that selects, tokenizes, and mixes data on the fly provides much more flexibility and eases experimentation.
-Utilizing a full-fledged DBMS for tracking data properties would burden ML engineers with complexities such as database administration, schema design, and performance tuning.
-We need a lightweight data plane that enables users to declaratively query and mix data based on arbitrary properties, independent of the filesystem structure, and to adjust this mixture dynamically.
+Running an offline data processing job with frameworks like Apache Beam or Spark to fully materialize the mixed training set for each training run makes it difficult to quickly iterate on different mixtures during the exploration phase and leads to data duplication, increasing storage costs.
+While databases offer a declarative query interface, using a fully-fledged DBMS to track data properties would burden ML engineers with database administration, schema design, and database performance tuning.
+Having a simple read-only SQLite table at each training client might seem like a straightforward solution, but it cannot dynamically adjust data mixtures and it does not interface with training frameworks like Torchtitan (Liang et al., 2024).
+We need a flexible data plane that lets users easily find and combine data based on any characteristic they choose, without being restricted by how files are organized.
+This system should also allow users to change the data mixture on the fly and work seamlessly with current training frameworks.
 | Mixtera | HF Datasets | WebDatasets | Mosaic Streaming | |||||||
@@
 | Checkpointing | yes | using TorchData | by replaying, not natively | yes |
-We present Mixtera, a data plane that can be deployed on top of existing LLM/VLM training data collections.
-It is a centralized, read-only layer and can be declaratively queried from training clients.
-As shown in Figure 2b, it reduces the amount of materialization and data scripting necessary for training jobs.
-While existing data loaders as outlined in Table 1 are limited by filesystem constraints, Mixtera supports arbitrary properties independent of the filesystem, and makes it easy for model engineers to experiment with different filter criteria, fixed learning curricula, and fully dynamic mixing algorithms that learn the mixture during training.
+We present Mixtera, a data plane that can be deployed on top of existing LLM/VLM training data collections stored in distributed filesystems or cloud data lakes.
 Mixtera follows a client-server model.
-For a training job, the server, which indexes and manages the samples, first statically filters out the relevant samples, and then continuously distributes chunks to the clients.
-Chunks are fixed-size collections of pointers to samples in files, and adhere to the current mixture.
-The clients then fetch the relevant samples and return the relevant data to the training loop.
-We design and implement Mixtera tailored to the needs for foundation model training, and make the following key contributions:
-Mixtera enables users to declaratively specify mixtures across properties independent of the filesystem structure and training frameworks by indexing all samples and their properties.
-Mixtera enables dynamically changing the data mixture and the properties determining the mixture during training.
+The user indexes all sample metadata at the server once.
+During a training job, the server statically filters out the relevant samples using SPJ-style predicates, and then continuously distributes chunks to the clients.
+Chunks are fixed-size collections of pointers to samples (Figure˜2b) in files that adhere to the current mixture.
+The clients fetch the relevant samples and return the relevant data to the training loop.
+By storing metadata, and deferring reads to the clients, Mixtera scales to real-world datasets.
+Chunks also enable dynamic data mixing and avoid re-materializing mixed data.
+This approach also improves the iteration speed for model developers.
+While existing data loaders, as outlined in Table˜1, rely on sampling from filesystem directories, Mixtera supports arbitrary properties independent of the filesystem, and makes it easy for model engineers to experiment with different filter criteria, fixed learning curricula (Allal et al., 2025; Xu et al., 2024b), and fully dynamic mixing algorithms that learn the mixture during training (Jiang et al., 2024; Chen et al., 2023, 2024a).
+We design and implement Mixtera tailored to the needs of foundation model training, and contribute the following:
+Mixtera indexes all samples and their properties, enabling users to declaratively specify mixtures across properties independent of the filesystem structure and training frameworks.
+Mixtera enables dynamically changing the data mixture and the properties used for mixing.
 It achieves this by generating and streaming chunks, i.e., fixed-size lists of pointers to samples following the current mixture.
-We show Mixtera does not bottleneck training and is versatile, enabling model accuracy improvements for text-only and text-vision models.
-As an example, we demonstrate how to implement the ADO dynamic data mixing algorithm in Mixtera and its positive impact on model accuracy.
+Mixtera’s data fetching scales to meet the ingestion throughput demands of large-scale training jobs.
+We run benchmarks spanning 256 GH200 superchips and show that Mixtera does not limit training throughput.
+As an example for how it enables model accuracy improvements, we demonstrate how to implement the ADO dynamic data mixing algorithm in Mixtera and its positive impact on model accuracy.
 2. Background
@@
 As of 2025, most such models are based on the Transformer architecture (Vaswani et al., 2017).
-They are trained on huge corpora of training data in a self-supervised manner to maximize the likelihood of predicting the tokens of a training sequence.
+They are trained on vast corpora of training data in a self-supervised manner to maximize the likelihood of predicting the tokens of a training sequence.
 Training phases.
 Training is structured into pre-training and post-training phases.
-In pre-training, we train a randomly initialized model on a general purpose data corpus (Section 2.1) to derive a base model.
+In pre-training, we train a randomly initialized model on a general-purpose data corpus (Section˜2.1) to derive a base model.
 In post-training, common steps include supervised finetuning (SFT) and alignment.
-The data used in SFT is human-selected for a specific task.
-In this paper, we focus on the pre-training use case.
+In this paper, we perform only pre-training experiments due to space constraints, but Mixtera can also be used for post-training.
 Distributed training.
 Training foundation models requires distributing computation across multiple GPUs.
-Training frameworks typically employ 3D parallelism (Ben-Nun and Hoefler, 2019; Hoefler et al., 2024), consisting of pipeline parallelism (PP), i.e., partioning the model layers between devices (Narayanan et al., 2021; Huang et al., 2019; Narayanan et al., 2019), tensor parallelism (TP), i.e., splitting individual tensor operations within layers across devices (Shazeer et al., 2018; Shoeybi et al., 2019), and data parallelism (DP), i.e., replication of the model across device groups.
+Training frameworks typically employ 3D parallelism (Ben-Nun and Hoefler, 2019; Hoefler et al., 2024), consisting of pipeline parallelism (PP), i.e., partitioning the model layers between devices (Narayanan et al., 2021; Huang et al., 2019; Narayanan et al., 2019), tensor parallelism (TP), i.e., splitting individual tensor operations within layers across devices (Shazeer et al., 2018; Shoeybi et al., 2019), and data parallelism (DP), i.e., replication of the model across device groups.
 PP and TP together are referred to as model parallelism.
@@
 2.1. Training Data and Data Mixing
-The data used for pre-training stems from data collections that include data from various sources, such as Wikipedia, Common Crawl dumps, or arXiv papers.
+Pre-training data stems from data collections that include samples from various sources (e.g., Wikipedia, Common Crawl dumps, or arXiv papers).
 Public examples of such collections include RedPajama (Weber et al., 2024), Dolma (Soldaini et al., 2024), and FineWeb (Penedo et al., 2024a).
-Besides the aggregation of data from different sources, data engineers typically clean the data, which typically involves deduplicating, filtering (e.g., the removal of personal identifiable information - PII), and applying classifiers to the data samples (e.g., to obtain a toxicity score for each sample).
-Longpre et al.
-(Longpre et al., 2024) and Penedo et al.
-(Penedo et al., 2024a) provide a great overview of the processing steps and their impact.
+Besides aggregating data from different sources, data engineers typically clean the data, which usually involves deduplicating, filtering (e.g., removing personal identifiable information), and applying classifiers to the data samples (e.g., to obtain a toxicity score for each sample) (Longpre et al., 2024; Penedo et al., 2024a).
 Data properties and mixtures.
-Each sample in the resulting collection has properties, such as its source (e.g., Wikipedia) or its language (e.g., English).
-Beyond the filtering operations, ML engineers need to define a data mixture.
-A mixture describes how the data is mixed based on their characteristics, i.e., we might train on 50 % data from Common Crawl and 50 % from movie subtitles.
+Each data sample has properties, such as its source (e.g., Wikipedia) or its language (e.g., English).
+ML engineers need to define a data mixture, which describes how the data is combined based on its characteristics, e.g., we can train on 50 % data from Common Crawl and 50 % from movie subtitles.
 The data can be combined based on multiple characteristics simultaneously.
 For instance, besides Common Crawl and movie subtitles, we might also use 80 % French and 20 % Italian data.
-The granularity on which the mixture is ensured depends on each training setup and there is no common agreement.
-For example, a mixture could be ensured within a batch, or across a window of several batches.
 Mixing algorithms.
@@
 We differentiate static mixtures, i.e., mixtures that remain constant over the entire training job, and dynamic mixtures, i.e., mixtures that change during the training job.
-In the last years, several algorithms for finding the best static mixture or how to adjust the mixture dynamically during training have been proposed.
-Algorithms such as DoReMi (Xie et al., 2024) or the data mixing laws (Ye et al., 2024) find the a static mixture via small proxy models.
+Recent research proposes several algorithms for finding the best static mixture or how to adjust the mixture dynamically during training.
+Algorithms such as DoReMi (Xie et al., 2024) or the data mixing laws (Ye et al., 2024) find a static mixture via small proxy models.
 Curriculum learning is an example of a pre-defined dynamic mixture.
@@
 The SmolLM2 model was trained on 4 stages of mixtures (Allal et al., 2025).
-Multilingual models are often first trained on English data, and then data from other languages is included (Richburg and Carpuat, 2024; Xu et al., 2024a).
-Beyond such pre-defined schedules, there is also work on adapting the data mixture to the model training dynamics, e.g., by increasing the weight of data domains which have high loss.
+Multilingual models are often first trained on English data, followed by samples from other languages (Richburg and Carpuat, 2024; Xu et al., 2024a).
+Beyond such pre-defined schedules, for text-only models, there is also work on adapting the data mixture to the model training dynamics, e.g., by increasing the weight of data domains that have high loss.
 Albalak et al.
-(Albalak et al., 2023) use a multi-armed bandit strategy to adjust the mixture.
+(Albalak et al., 2023) model mixture components as arms of a multi-armed bandit.
 Skill-it orders “skills” based on model feedback (Chen et al., 2023).
@@
 PiKE relies on gradient interactions (Li et al., 2025).
-In this paper, we use Adaptive Data Optimization (ADO) (Jiang et al., 2024) as an example of a dynamic mixing algorithms.
+In this paper, we use Adaptive Data Optimization (ADO) (Jiang et al., 2024) as an example of a dynamic mixing algorithm.
 2.1.1. Adaptive Data Optimization
 Adaptive Data Optimization (ADO) is a dynamic mixing algorithm that adjusts the data mixture during training based on the model’s learning progress on each domain (Jiang et al., 2024).
-The key idea is to prioritize domains where the model shows rapid improvement while considering how much each domain contributes to its own progress.
+The key idea is to prioritize domains where the model shows rapid improvement while considering how much each domain benefits from its own samples.
 ADO uses neural scaling laws to model how the loss of each domain decreases with the number of training samples .
@@
 Additionally, ADO enforces a minimum sampling probability for each domain.
-For more details, we refer to Jiang et al.
-(Jiang et al., 2024).
 3. Current Challenges
 We identify three challenges in the status quo of training data management with current open-source infrastructure.
-Challenge 1: High engineering effort for data preparation.
-The current approach to preparing training data sets involves many manual offline steps with general-purpose data processing and scripting frameworks (Figure 2a).
-For the offline cleaning step (Section 2.1), ML engineers typically leverage data processing frameworks like Spark (Zaharia et al., 2016), Beam (Akidau et al., 2015), Data-Juicer (Chen et al., 2024b), or datatrove (Penedo et al., 2024b).
-The subsequent data mixing can happen offline or online.
-If data is mixed offline, engineers write ad hoc scripts which create a new mixed copy of the cleaned dataset for each training run.
-This makes it difficult to get a quick sense of how a data mixture will impact model training when exploring different mixing policies.
-Some existing data loaders support online data mixing.
-However, they only implement it across directories, i.e., they assign a weight to each directory and sample data according to the weights.
-Hence, the directories need to reflect the property we want to mix on, e.g., one directory per language or data source.
-This inflexible approach neither supports switching the property we mix on nor supports specifying hierarchical mixtures across arbitrary properties, e.g., specify a proportion between Wikipedia and movie subtitles, and between English and German (c.f.
-Table 1).
-Challenge 2: Inefficient data management on filesystems.
-Training data is typically stored and managed as files without a proper data management system, which can lead to storage overhead, performance bottlenecks, and consistency issues.
-Both the source data and mixed training set are typically stored in formats such as jsonl or parquet files in a shared, distributed filesystem.
-Existing data loaders just wrap around the filesystem and are therefore limited by its constraints (Table 1).
-Filesystems are commonly used because current database management systems are not natively built for foundation model training data (Wang et al., 2023).
-Using a proper DBMS for such data would require ML developers to define table schemas and to orchestrate dataflow from the DBMS to the model training, in addition to administration overhead to operate the DBMS itself.
-This situation is problematic, as (i) both the metadata and actual payloads are commonly duplicated across training jobs, (ii) filesystems do not provide an efficient interfaces to query data, (iii) there is no native way of tracking which model was trained on which data if it is accessed via general-purpose filesystem calls.
-Some frameworks like Megatron (Shoeybi et al., 2019; Narayanan et al., 2021) even require pre-tokenized data, which leads to duplication of the even bigger tokenized data files.
-Challenge 3: Rapidly evolving research.
-How to train the best model on a given dataset is an active area of research, with many new techniques such as dynamic mixing emerging.
-Offline preparation of the mixture or online mixing based on fixed directory weights does not support dynamic mixture at all.
-Even for researchers who are familiar with the latest mixing techniques, implementing modern mixing algorithms in a training pipeline is a painful, tedious, and error-prone task.
-The codebases for mixing algorithms are often tailored directly to the training framework, as well as the data collection and properties used in the respective papers.
-This hinders the adoption of new methods and makes researching, reproducing, and comparing different strategies difficult.
+Challenge 1: Today’s training data storage systems lack expressive, declarative APIs for data mixing.
+Training data is typically stored and managed as files on distributed filesystems or objects on cloud storage.
+As these systems are not natively built for foundation model training data (Wang et al., 2023), they lack operators for data selection, mixing, and ingestion into a training framework.
+This also complicates lineage tracking, as there is no native way of tracking which model was trained on which data when the data is accessed via general-purpose filesystem calls.
+Challenge 2: Preparing and materializing data mixtures with offline preprocessing limits flexibility and increases storage costs.
+The current approach to preparing training data involves numerous manual offline steps with general-purpose data processing and scripting frameworks (Figure˜2a).
+For the offline cleaning step (Section˜2.1), ML engineers typically leverage data processing frameworks like Spark (Zaharia et al., 2016), Beam (Akidau et al., 2015), Data-Juicer (Chen et al., 2024b), or datatrove (Penedo et al., 2024b).
+Subsequent data mixing can happen offline or online.
+Current online data loaders sample data from directories that reflect the mixing property, e.g., one directory per source dataset (Table˜1), and therefore do not support switching the property we mix on nor specifying hierarchical mixtures across arbitrary properties.
+Due to the limited functionality of current online solutions, engineers often write ad hoc offline mixing scripts that create a new mixed copy of the cleaned dataset for each training run, which can drastically increase storage usage and cost.
+As some frameworks like Megatron (Shoeybi et al., 2019; Narayanan et al., 2021) even require pre-tokenized data, this leads to duplication of the even bigger tokenized data files.
+Offline mixing also drastically increases the iteration time and makes it difficult to get a quick sense of how a mixture will impact model training when exploring different mixing policies.
+Challenge 3: Lack of support for dynamic mixtures during training.
+How to train the best model on a given dataset is an active research area, with dynamic mixing emerging as a prominent new technique.
+Offline preparation of the mixture or other approaches such as online mixing based on fixed directory weights or using a vanilla DBMS without additional infrastructure does not support dynamic mixture at all.
+Additionally, even for researchers who are familiar with the latest mixing techniques, implementing modern mixing algorithms in a training pipeline is a painful, tedious, and error-prone task, as the codebases for mixing algorithms are often tailored directly to the training framework, as well as the data collection and properties used in the respective papers.
+This hinders the adoption of dynamic mixing algorithms and makes researching, reproducing, and comparing difficult.
 4. Mixtera’s Design
-We propose to address these challenges for training data management by building Mixtera, a foundation model training data plane. We derive the following design goals for such a system.
+We address these challenges for training data management by building Mixtera, a foundation model training data plane. We derive the following design goals for such a system.
 Goal G1: The system should implement a centralized data layer that users can conveniently and declaratively query to mix data across arbitrary properties, independent of the filesystem structure.
-Goal G2: The system needs to be lightweight, i.e., not require many components to set up and be easily integratable into existing training setups.
-Goal G3: While being user-friendly and flexible, the system needs to ensure high-throughput and determinism (reproducibility).
-Goal G4: The system must support adjusting the mixture dynamically during the training.
-4.1. System Overview and Design
-TODO!
-Inspired by the Lakehouse architecture (Armbrust et al., 2021), we design Mixtera as a read-only layer that can be deployed on top of existing training data collections, which are typically stored in a distributed filesystem.
-Figure 3 shows the architecture of the system.
-Mixtera manages a centralized database of metadata (i.e., properties such as language or source dataset) of all training data samples (G1).
-A sample can be a piece of text (for LLM training) or a text-image pair (for VLM training).
-Mixtera assigns every sample a unique ID and properties instead of treating the training data as a blob of homogeneous, contiguous data.
-It allows users to declaratively query the relevant samples for a training.
-To be lightweight (G2), Mixtera does not reorganize or modify the data files on disk.
-For model training, it provides a standard iterator that can be used, e.g., in conjunction with a torch.DataLoader.
-Mixtera is agnostic to the model training framework, supports training interruptions using checkpoints, and ensures determinism (G3) through careful shuffling, i.e., for identical queries, Mixtera always provides data in an identical order, which is important for reproducibility as well as for debugging issues like loss spikes (Zhuang et al., 2022; Cooper et al., 2022; Qian et al., 2021; Karamcheti et al., 2021; Team, 2025).
+Goal G2: The system needs to be lightweight, i.e., easily integrate into existing training deployments without requiring setting up many components.
+Goal G3: The system needs to ensure high-throughput, determinism, and reproducibility, while being user-friendly and flexible.
+Goal G4: The system must support adjusting the mixture dynamically during training.
+4.1. Data Model
+Mixtera operates on training datasets of the following structure:
+Sample.
+A sample is the atomic unit of training data, as supplied by the user.
+In LLM training, the sample unit is typically a text string or a pre-tokenized sequence; VLM training uses multimodal samples (e.g., text-image pairs).
+Sample granularity varies in practice, ranging from short texts to full documents.
+Property.
+A property is a named attribute of a sample, with a domain of possible values.
+Properties characterize samples and can be single-valued (each sample has exactly one value, e.g., ), or multi-valued (each sample may have multiple values, e.g., ).
+Data Collection.
+A data collection is a tuple where is a finite set of samples, is a set of properties, and maps each (sample, property) pair to a subset of values from that property’s domain.
+We also refer to it as the sample metadata.
+Mixtera defines mixtures fully independent of the filesystem–it is fully based on (logical) properties.
+Mixture Key. A mixture key is a partial mapping from a subset of properties to sets of values, i.e., , with . For mixture keys and , we say matches , if f.a. , . This relation is non-symmetric.
+Component Key. A sample induces a component key , s.t. .
+Data Mixture. A data mixture is a mapping from mixture keys to target proportions: with .
+4.2. System Overview
+Mixtera is a read-only layer that can be deployed on top of existing sets of samples , which are typically stored in a distributed filesystem or cloud object store.
+Figure˜3 shows the client-server architecture of the system.
+The server runs on one node, and each training node runs client instances.
+Mixtera manages a centralized database which stores the mapping , i.e., the sample metadata (G1).
+This database needs to be populated by scanning all samples once before it can be used for training.
+Since the metadata typically consists of integers and floats, it is much smaller than the actual samples.
+Mixtera defers the reading of actual sample payloads to clients.
+This enables to scale to large-scale training datasets.
+Mixtera assigns every sample a unique ID.
+It allows model developers to declaratively query the relevant samples for a training job.
+To remain lightweight (G2), Mixtera does not reorganize or modify the data files on disk.
+It provides a standard iterator for model training that can be used in conjunction with loaders like torch.DataLoader.
+Mixtera is agnostic to the training framework, supports training interruptions using checkpoints, and ensures determinism (G3) through careful shuffling, i.e., for identical queries, Mixtera always provides data in an identical order, which is important for reproducibility and debugging issues like loss spikes (Zhuang et al., 2022; Cooper et al., 2022; Qian et al., 2021; Karamcheti et al., 2021; Team, 2025).
 It supports adjusting the mixture during training (G4) by transferring chunks (lists of pointers to samples) whose mixture can change over time.
-[ frame=lines, framesep=2mm, baselinestretch=0.8, bgcolor=LightGray, fontsize=, linenos ] python client = MixteraClient("127.0.0.1", 8080) job_id = "test_job" query = Query.for_job(job_id).select(("license","==","CC")) mixture = StaticMixture( MixtureKey("language": ["JavaScript"]): 0.7, MixtureKey("language": ["HTML"]): 0.3 , chunk_size=1024) qea = QueryExecutionArgs(mixture=mixture, num_workers=4, dp_groups=1, nodes_per_group=1) rsa = ResultStreamingArgs(node_id=0, dp_group_id=0, job_id=job_id) ds = MixteraTorchDataset(client, query, rea, rsa) dl = torch.utils.data.DataLoader(ds, batch_size=1024, num_workers=4)
-[…]
-Query interface.
-In Figure 4, we show an example of a query that statically selects only Creative Commons data, and then mixes HTML and JavaScript data in a 70:30 ratio during training.
-Mixtera’s implementation takes care of executing the query and obtaining the samples without needing to worry about correctness, even in distributed training.
-The user only needs to provide the ID of the node and its data parallel group, which is obtained from the training framework.
-Mixtera currently allows to express static filter operations on properties, and static as well as dynamic mixtures across all properties (G4).
-For more details on supported mixture types, we refer to Section 5.2.2.
-Dataflow.
-Mixtera follows a server-client model.
-As shown in Figure 3, the server runs on a node and each training node runs client instances.
-Users first register samples at the server to populate the metadata database.
-They then can submit queries.
-A query is executed at the server in two phases.
-First, Mixtera applies static filters from the query (e.g., English-only) to obtain all samples eligible for training.
-Second, during training, the server distributes chunks of that query result to the client(s), which specify which samples to train on.
-The server ensures that the chunks are distributed correctly, i.e., tensor- and pipeline parallel stages receive the same input data.
-The server generates chunks according to the current mixture, i.e., it iteratively takes samples from the query result generated in the first phase of query execution, such that data in the chunk follow the current mixture.
-As an iterable data loader, Mixtera faces the challenges of determinism and checkpointing.
-We address this by shuffling based on the query and support to load/store the query state.
+Data types.
+Mixtera supports diverse sample types (e.g., pre-tokenized text in binary files, strings in jsonl/ parquet, text–image pairs in webdatasets).
+This covers many relevant use cases.
+Even modern foundation models for tabular data treat tables as strings (Bodensohn et al., 2025).
+The sample granularity is inherent to the ingested dataset.
+As Mixtera is a read‑only layer, it adopts whatever granularity the data already uses.
+While some features such as tokenization or image decoding are modality-specific, the system is fundamentally agnostic to the sample unit.
 Chunks.
-Mixtera does not store the sample payloads, but rather only the metadata of each sample.
-During training, the nodes receive chunks.
-A chunk is a fixed-size collection of pointers to samples in files.
-They tell the client which samples which file to load and train on (e.g., sample 10 in file wikipedia.jsonl.zst).
-The files can be local, in a distributed filesystem, or cloud storage.
-Storing and distributing pointers to samples instead has several advantages.
-First, users can store data in their locations of choice (e.g., an object store, or an distributed filesystem).
+Chunks are Mixtera’s core abstraction for scheduling and mixture enforcement.
+Abstractly, a chunk is a collection of samples of fixed size , and the set conforms to the current mixture .
+That is, for each mixture key , the proportion of samples in matching is approximately .
+In implementation, Mixtera does not materialize or transfer the sample payloads .
+Instead, chunks contain only pointers to samples: metadata specifying which samples in which files to load (e.g., “samples 100-150 in wikipedia.jsonl.zst”).
+Deferred reading.
+Storing only metadata and distributing pointers to samples (as opposed to ingesting the actual sample payloads into the system) has several advantages.
+First, users can store data in their locations of choice (e.g., an object store, or a distributed filesystem).
 Chunks are independent of the filesystem structure (G1).
@@
 The server only creates chunks and each client fetches the data they need.
-Fourth, we avoid a lock-in effect on Mixtera and allow for easy adoption on existing data collections (G2).
-Last, we natively support other modalities like images, which would not be straightforward to ingest at scale into a database.
-Mixtures.
-Users can use any subset of sample properties to define a mixture.
-Users can even change the properties that they mix on during training with dynamic mixture schedules.
-To handle this, Mixtera implements a MixtureKey abstraction that allows users to describe which samples they want to use.
-The keys also implement flexible property matching to define which samples to use during chunk generation (Section 5.2.2).
+Fourth, we avoid a lock-in effect and allow for easy adoption on existing data collections (G2).
+Last, we natively support other modalities (whereas image or video payloads would not be straightforward to ingest at scale into a database).
+Query interface.
+In Figure˜4, we show an example query that statically selects only Creative Commons data, and then mixes HTML and JavaScript data in a 70:30 ratio during training.
+Mixtera takes care of executing the query and obtaining the samples without needing to worry about correctness, even in distributed training.
+The user only needs to provide the ID of the node and its data parallel group, which is obtained from the training framework.
+Mixtera allows expressing static filter operations on properties via SPJ-style predicates, and static as well as dynamic mixtures across all properties (G4).
+The static filter defines the ground set of all potential data (e.g., only CC), while the mixture describes the proportions of how the data from this set is mixed during training (JavaScript:HTML ratio).
+The MixtureKey class operationalizes its formal definition (Section˜4.1) in a user-friendly Python API.
+Executing a query.
+Before submitting a query, users catalog all samples at the server.
+This involves defining the data schema as a Python class, which allows Mixtera workers to scan all samples and populate the database.
+Users then can submit queries.
+A query is executed at the server in two phases.
+First, Mixtera applies static filters from the query (e.g., English-only) to obtain all samples eligible for training (QueryResult).
+Second, during training, the server distributes chunks of that query result to the client(s), which specify which samples to train on.
+The server ensures that the chunks are distributed correctly, i.e., tensor- and pipeline parallel stages receive the same input data.
+The server generates chunks according to the current mixture, i.e., it iteratively takes samples from the query result such the chunk abides by the current mixture.
+As an iterable data loader, Mixtera faces the challenges of determinism and checkpointing.
+We address this by shuffling based on the query and support to load/store the query state.
 High-throughput data fetching.
-The challenge of using chunks is that we need to handle suboptimal data layout.
-Files may have arbitrary property distributions, or might follow a partial structure (e.g., a file only contains Wikipedia data, but the languages are distributed randomly).
-We cannot re-organize data within the existing files.
-Formats like jsonl were not built with random access in mind, but chunks force clients to load individual samples from files.
-To avoid data stalls (G3), we implement chunk generation to use subsequent sample ranges in files if the samples have the same properties and engineer Mixtera’s implementation to fetch those ranges as efficiently as possible (Section 5.3).
-Open source.
+The challenge with not re-organizing the user’s data files is that Mixtera needs to handle suboptimal data layout.
+Files may have arbitrary distributions of data properties (e.g., even if a file only contains Wikipedia data, data in different languages might be distributed randomly across files).
+Formats like jsonl were not built with random access in mind, yet chunks force clients to load individual samples from files.
+To avoid data stalls (G3), Mixtera reads subsequent samples in files if the samples follow the same properties (e.g., same language).
+This is achieved through intervals identifying consecutive samples in file that share the same component key .
+A chunk is thus represented as a collection of intervals whose total size equals and whose union satisfies the mixture.
+We use nested CTEs in DuckDB to find these intervals as fast as possible (Section˜5.3).
+Users are not required to use sequential file formats.
+Alternative design considerations.
+A metadata SQLite database at each client might seem like a simple alternative.
+While DBs can federate and query external sources with SPJ-style predicates, we still need an interface between the storage, query engine, and training framework, which can turn query results into input tensors (G1, G2).
+Furthermore, using a database at each client would not enable dynamic mixing algorithms (G4).
+Mixtera serves as the interface that connects a metadata database to the training framework, with a full feature set for dynamic model training.
+It also alleviates ML engineers from manually maintaining the database.
+We do not find centralized metadata management to be a bottleneck even for large-scale training jobs, because the work at the service is minimal (occasional chunk creation and distribution).
+Mixtera maintains high training throughput for large training jobs (Section˜6.3).
+Industry frameworks like Google’s Pathways (Barham et al., 2022) also rely on a single Python process to coordinate large-scale training.
+A current limitation of Mixtera’s design is that it does not enable incorporating new data into an existing query.
+However, this (i) affects nearly all dataloaders as they all build some internal state, and (ii) and can in all cases be dealt with by executing a new query, e.g., in continual pretraining from a model checkpoint.
+Open source ecosystem.
 Mixtera comes as a Python package that provides the entry point for the server and abstractions for the client.
-Its codebase, consisting of approximately 11 k lines of Python and C++ (excluding tests), is open-source111Available at https://github.com/eth-easl/mixtera..
+The codebase, consisting of approximately 11 k lines of Python and C++ (excluding tests), is open-source111Available at https://github.com/eth-easl/mixtera..
 It is rigorously tested with a full set of unit and integration tests.
-We are continuously working on enhancing the system and welcome contributions.
+We are continuing to add features and welcome contributions.
 5. Implementation
-We explain how sample metadata is managed (Section 5.1) at the server, how it executes queries and creates chunks (Section 5.2), how those chunks are parsed at the client (Section 5.3), and give details on Mixtera’s integration into training frameworks (Section 5.4).
-5.1. Sample Management
-Mixtera manages samples in a MixteraDataCollection (MDC).
-It keeps track of all samples and their properties.
-The MDC leverages DuckDB (Raasveldt and Mühleisen, 2019) as the robust and flexible metadata management system underlying it.
-Every registered sample has properties according to a schema defined by a MetadataParser, e.g., language.
-Representing the structure of training data, Mixtera initializes the database with three tables for source datasets, files, and samples.
-The source dataset can also be ignored and stored as a schema property instead, to support collections such as The Pile (Gao et al., 2020) where data from different datasets is contained within the same files.
-MetadataParsers.
-A MetadataParser is a class that defines a schema for a data collection and, given a sample, extracts and returns the properties according to the schema.
-A schema is a list of properties which have a type (e.g., string or enum), a nullable field, and a multiple field, describing whether a single sample can take multiple values for this properties, e.g., multiple languages.
-Mixtera maps this Python schema to a proper database schema, comes with a set of pre-defined parsers for common datasets, and enables users to define custom parsers.
-For example, there is a parser for The Pile (Gao et al., 2020).
-Every sample has the property pile_set_name describing the source dataset.
-This is a non-nullable, non-multiple, enum property, since we know the limited set of possible values.
-Data types.
-Based on the parser, Mixtera adjusts the DuckDB sample table.
-Properties marked as multiple get mapped to DuckDB’s list type, e.g., a multiple string property gets mapped to VARCHAR[].
-Enum properties are appropriately inserted into DuckDB.
-Using enum instead of strings for properties where all values are known beforehand speeds up query execution in DuckDB.
-Insertion.
-When inserting (registering) data, the MDC accumulates all relevant files, and prepares the DuckDB table schema.
-Then, all files are inserted into the files table, and a pool of workers in parallel processes all files using the MetadataParser to extract the sample metadata.
-We then aggregrate the worker results and insert it into the database, as DuckDB does not support insertion from multiple Python process workers.
-We found that converting the worker results to columnar pyarrow in-memory tables and then inserting into the samples table has the highest throughput.
+We explain how Mixtera ingests sample metadata (Section˜5.1), how it executes queries and creates chunks (Section˜5.2), how those chunks are parsed at the client (Section˜5.3), and describe Mixtera’s integration into training frameworks (Section˜5.4).
+5.1. Metadata Insertion
+Mixtera manages the metadata in an MixteraDataCollection (MDC) which uses DuckDB (Raasveldt and Mühleisen, 2019) as the underlying DBMS to efficiently store and query sample properties.
+The implementation itself is agnostic to the DBMS.
+While DuckDB’s pluggable extension architecture (or similar approaches like PostgreSQL’s Foreign Data Wrappers or SQLite’s Virtual Tables) would allow to query the sample files directly, to defer reading (Section˜4.2) to the training nodes, the MDC serves as a metadata index.
+Initial ingestion.
+To populate the MDC, users need to define a MetadataParser.
+A MetadataParser operationalizes the property set and extraction of for each sample .
+It defines the schema specifying and domains for each property.
+In Python, the schema is a list of properties which have a type (e.g., string or enum), a nullable field, and a multiple field, describing whether a single sample can take multiple values for this properties (e.g., several languages).
+Mixtera adjusts the underlying database table by mapping the Python schema to a proper database schema.
+The system comes with a set of pre-defined parsers for common datasets and enables users to define custom parsers.
+File scanning.
+Metadata ingestion is a one-time preprocessing step.
+The MDC first accumulates all data files and prepares the database schema.
+Mixtera then parallelizes metadata extraction using a worker pool.
+Workers process files in batches, where each worker sequentially reads samples from its assigned files and applies the MetadataParser to extract property values (e.g., parsing JSON fields).
+Since DuckDB does not support concurrent insertions from multiple processes, Mixtera aggregates worker results in the main process.
+To optimize insertion throughput, we convert the collected metadata to columnar PyArrow in-memory tables before bulk insertion into DuckDB.
+Importantly, metadata extraction is decoupled from training.
+Users must define properties upfront, but can later add new properties by re-scanning the dataset with an updated parser without affecting existing metadata.
 5.2. Server-Side Query Execution
-After receiving a query from the client, the server executes it in two phases.
-The first phase is performed via the MDC DuckDB-abstraction and applies static filters to identify all potentially relevant samples, e.g., all non-English samples are filtered out, and groups consecutive samples into intervals.
-The second phase constructs a specialized data structure called ChunkerIndex that enables efficient, mixture-aware chunk generation.
+After registering data, users can execute queries.
+When a client sends a query, the server executes it in two phases.
+The first phase is performed via the MDC and applies static filters to identify all relevant samples, and groups consecutive samples into intervals.
+The second phase constructs a data structure called ChunkerIndex that enables efficient, mixture-aware chunk generation.
 5.2.1. SQL generation and interval detection
 After receiving an object representation of the query (c.f.
-Figure 4), similar to ORM frameworks like sqlalchemy, the Mixtera server generates a base SQL query from this object.
+Figure˜4), similar to ORM frameworks like sqlalchemy, the Mixtera server generates a base SQL query from this object.
 This query returns a table in which each row represents a sample that the user is interested in.
@@
 Even though samples 4-5 (file 1) and 1 (file 2) share the same properties (Python, Apache), they are in different files and thus form separate intervals.
-The primary key is formed by the sample and file ID together.
+The primary key is formed by the sample and file ID.
 Mixtera constructs a SQL query that processes the data in multiple stages:
-First, Mixtera establishes a Common Table Expression (CTE) named base_data that contains the filtered samples: {minted}[frame=lines,framesep=2mm,baselinestretch=0.8,bgcolor=LightGray,fontsize=]sql WITH base_data AS ( – Our generated base filtering query here, e.g., SELECT * FROM samples WHERE license = ’MIT’),
+First, Mixtera establishes a Common Table Expression (CTE) named base_data that contains the filtered samples:
+WITH base_data AS (– Our generated base filtering query here, e.g.,SELECT * FROM samples WHERE license = ’MIT’), -
 Next, Mixtera identifies breaks in the sample sequence using window functions.
-The grouped_samples CTE calculates the difference between consecutive sample IDs within groups sharing the same properties: {minted}[frame=lines,framesep=2mm,baselinestretch=0.8,bgcolor=LightGray,fontsize=]sql grouped_samples AS ( SELECT *, sample_id - LAG(sample_id, 1, sample_id) OVER (PARTITION BY file_id, lang, license ORDER BY sample_id) AS diff FROM base_data), Here, a diff value of 1 indicates consecutive samples, while any other value indicates a break in the sequence.
-The intervals CTE then groups the sequences into intervals: {minted}[frame=lines,framesep=2mm,baselinestretch=0.8,bgcolor=LightGray,fontsize=]sql intervals AS ( SELECT file_id, lang, license, SUM(CASE WHEN diff != 1 THEN 1 ELSE 0 END) OVER (PARTITION BY file_id, lang, license ORDER BY sample_id) AS group_id, MIN(sample_id) as int_strt, MAX(sample_id)+1 as int_end FROM grouped_samples GROUP BY file_id, lang, license, diff, sample_id) The group_id is incremented when there is a break in the sequence, creating unique identifiers for each interval.
-Finally, Mixtera aggregates the results to get the final intervals: {minted}[frame=lines,framesep=2mm,baselinestretch=0.8,bgcolor=LightGray,fontsize=]sql SELECT file_id, lang, license, group_id, MIN(int_strt) as interval_start, MAX(int_end) as interval_end FROM intervals GROUP BY file_id, lang, license, group_id ORDER BY file_id, interval_start;
-While we initially implemented the interval aggregation within Python, letting DuckDB optimize and parallelize the calculation of the interval is more efficient.
-This optimization can only improve I/O if samples within files are clustered by properties and not randomly distributed.
-For example, if a user has a file structure based on the source dataset and mixes based on that, Mixtera can heavily leverage the range optimization and just read, e.g., all Wikipedia samples from the sample file.
-If the user decides to now mix on language as well, there might only be local clusters within the files.
-Since Mixtera is read-only by design, it cannot re-shuffle data.
+The grouped_samples CTE calculates the difference between consecutive sample IDs within groups sharing the same properties:
+grouped_samples AS (SELECT *, sample_id - LAG(sample_id, 1, sample_id)OVER (PARTITION BY file_id, lang, licenseORDER BY sample_id) AS diffFROM base_data),Here, a diff value of 1 indicates consecutive samples, while any other value indicates a break in the sequence.
+The intervals CTE then groups the sequences into intervals:
+intervals AS (SELECT file_id, lang, license,SUM(CASE WHEN diff != 1 THEN 1 ELSE 0 END)OVER (PARTITION BY file_id, lang, licenseORDER BY sample_id) AS group_id,MIN(sample_id) as int_strt, MAX(sample_id)+1 as int_endFROM grouped_samplesGROUP BY file_id, lang, license, diff, sample_id)The group_id is incremented when there is a break in the sequence, creating unique identifiers for each interval.
+Finally, Mixtera aggregates the results to get the final intervals:
+SELECT file_id, lang, license, group_id,MIN(int_strt) as interval_start, MAX(int_end) as interval_endFROM intervalsGROUP BY file_id, lang, license, group_idORDER BY file_id, interval_start;
+Using intervals of samples can only improve I/O if samples within files are clustered by properties and not randomly distributed. Since Mixtera is read-only by design, it does not re-shuffle data.
 5.2.2. Chunk generation
-The chunk generation algorithm is based on the ChunkerIndex data structure, which organizes sample ranges by their properties.
-This data structure is generated based on the output from the interval query.
-Before explaining its construction, we first introduce the key abstraction that enables flexible property matching: MixtureKey.
+After obtaining the query result with all relevant intervals, the server next runs the chunk generation algorithm.
+This algorithm is based on the ChunkerIndex data structure, which organizes sample ranges by their properties.
+We re-visit the MixtureKey concept from an implementation perspective.
 MixtureKey abstraction.
-A MixtureKey represents a set of properties and their values, e.g., language being English and license being MIT.
-A property in a key can have multiple values.
-A key matches another if their lists of values intersect for all properties they share.
-This matching is crucial because the resulting interval table from DuckDB contains the full cross-product of all properties—a sample might have values for language, license, size, topic, and more—while a mixture specification typically considers only a subset of these properties.
-For example, a mixture might specify language:English to select English text regardless of license, while the underlying samples are stored with full property information (e.g., language:English,German;license:CC and language:English;license:MIT).
-The flexible matching enables finding all samples that satisfy certain properties while ignoring irrelevant attributes they might have.
+A MixtureKey represents a set of properties and their values.
+The class implements the formal definition from Section˜4.1: a partial mapping for some .
+The matching relation enables flexible querying: matches if .
+This matching is crucial as the resulting interval table from DuckDB contains the full cross-product of all properties—a sample might have values for language, license, size, topic, and more—while a mixture specification may consider only a subset of these properties, as we discuss in the next paragraph.
 It also allows us to define mixtures on multiple properties with multiple values, instead of being limited to a single property (c.f.
-Section 3).
+Section˜3).
 To ensure deterministic behavior, we implement a total ordering over keys based on the number of properties, property names, and their values.
-We sometimes refer to a specific MixtureKey as a domain, e.g., the key for language:English defines the domain for English samples.
+We sometimes refer to a specific MixtureKey as a domain, e.g., the key for lang:English defines the domain of English samples.
 The ChunkerIndex.
-The ChunkerIndex is a nested mapping from MixtureKeys to sample locations.
-We refer to keys in the ChunkerIndex as component keys.
-For each unique combination of all available properties and values, the index maps to dataset IDs, which in turn map to file IDs and finally to sorted lists of intervals of samples which share the properties and values as specified in the MixtureKey.
-While the index maintains the complete property information of samples, the matching semantics allows us to efficiently query arbitrary subsets of these properties.
-Consider a simplified example where we represent MixtureKeys as strings.
-A fragment of the ChunkerIndex might look like: {minted}[frame=lines,framesep=2mm,baselinestretch=0.8,bgcolor=LightGray,fontsize=]python "language:JavaScript,HTML;license:MIT": ds_1: file_1: [(1,4), (10,15)], # half (right) open ranges file_2: [(1,2)] , "language:Python;license:Apache": ds_1: file_1: [(1,2)]
+Recall from Section˜4.1 that each sample induces a component key defined f.a.
+as .
+The ChunkerIndex organizes intervals by these component keys, i.e., , where is the filtered subset from the query, and intervals group consecutive samples sharing the same .
+While the index maintains the complete property information of samples, it enables efficient sample lookup: given a mixture key , we identify all component keys where matches , and retrieve their associated intervals.
+Consider a simplified example with MixtureKeys as strings.
+A fragment of the ChunkerIndex might look like:
 In this example, a query for language:JavaScript would match the first key despite it having the additional license property and two assigned languages.
@@
 Using multiprocessing to circumvent the GIL would require expensive pickling of nested dictionaries.
-Our C++ implementation uses multithreading and only acquires the GIL at the end when creating the final Python result object.
+Our C++ implementation uses multithreading and only acquires the GIL at the end.
 Each C++ worker thread maintains a local index for a subset of the data.
@@
 Chunk generation.
-A mixture in Mixtera is represented as a mapping from MixtureKeys to their target proportions within a chunk.
-Given a chunk size that describes how many samples each chunk should contain, these proportions are converted to absolute counts following the largest remainders quota method.
-Users can set a mixture to be strict, requiring exact proportions, or best-effort, allowing to continue to generate chunks even if the mixture cannot be exactly fulfilled.
-The chunk generation algorithm (Algorithm 1) iteratively constructs chunks that conform to the current mixture.
+As defined in Section˜4.2, a chunk is abstractly a collection of samples conforming to mixture .
+In implementation, contains only interval pointers rather than materializing samples.
+Given and chunk size , Algorithm˜1 constructs such that for each mixture key , the number of samples with component keys matching approximates .
+Users can set a mixture to be strict, requiring exact proportions, or best-effort (continue to generate chunks even if the mixture cannot be exactly fulfilled).
+The algorithm iterates through each , identifies all matching component keys in the ChunkerIndex, and extracts intervals until the target count is reached.
 This algorithm supports dynamic mixture, as the mixture can be changed between chunks.
-Similar to the ChunkerIndex, a chunk is also structured as a hierarchical dict with keys and files pointing to lists of ranges.
-This chunk can be parsed on a training node by loading the sample ranges from the files (Section 5.3).
-While the ChunkerIndex uses all properties/values as MixtureKeys (component keys), a chunk’s keys are identical to the mixture’s keys, which might only contain a subset of properties.
 For each key in the mixture, the algorithm keeps track of how many samples we still need to put into the chunk that is currently being generated (remaining_sizes).
@@
 A call to obtain samples for a component key can return fewer samples than requested, e.g., if we are looking for JavaScript data and we need 5 samples, but we only have 3 JavaScript/MIT licensed samples, the according component key can only return 3 samples.
-Requesting samples is implemented as requesting ranges (from potentially multiple files) such that the overall number of samples in the returned list of ranges is .
-The lists of ranges per file are merged into the existing sorted list of ranges.
-Mixtera’s implementation uses Python generators that yield ranges of samples and accept the number of samples needed as input through the Python generator’s send mechanism.
-This hides the complexity of the take samples operation.
-It also allows for efficient, stateful iteration over the available ranges while maintaining control over how many samples are returned in each request.
-There is one generator per component key that returns the ranges containing samples based on the ChunkerIndex but ensures ranges are split such that never too much data is returned.
-If after traversing all component keys we did not find sufficient samples for a key in the mixture, in strict mode, chunk generation fails.
+Requesting samples is implemented as requesting intervals (from potentially multiple files) such that the overall number of samples in the returned list of intervals is .
+The lists of intervals per file are merged into the existing sorted list of intervals.
+If, after traversing all component keys, we did not find sufficient samples for a key in the mixture, in strict mode, chunk generation fails.
 In best-effort mode, the algorithm redistributes any unfulfilled counts to the remaining mixture components proportionally to their original ratios.
@@
 To avoid infinite loops, we only distribute samples to keys on which we were able to find any samples in the last iteration.
-Mixtera’s implementation of this algorithm ensures determinism because (1) the mixture’s keys are processed in a consistent order and (2) when multiple component keys match a mixture key, they are considered in a deterministic order based on a seeded shuffle of all possible keys.
-This approach ensures that identical queries with identical mixtures always produce identical chunks, important for training debugging and reproducibility (Zhuang et al., 2022; Cooper et al., 2022; Qian et al., 2021; Karamcheti et al., 2021; Team, 2025).
-Mixtera also supports an alternative, arbitrary chunk generation algorithm.
-It does not require any specified mixture and builds up chunks by iterating over the component keys, moving on to the next one when one key is depleted.
-This can be useful if the data is completely property-free and guarantees maximum throughput as all ranges are consecutive.
+This redistribution mechanism can be enabled or disabled, allowing users to either prioritize strict mixture fidelity (which may stop training when samples are exhausted) or training as long as possible with approximated proportions.
+Implementation details.
+The take samples operation (line 9) is implemented using Python generators that yield ranges of samples and accept the number of samples needed as input through the generator’s send mechanism.
+This hides the complexity of range management and allows for efficient, stateful iteration over available ranges while maintaining control over sample counts.
+There is one generator per component key that returns ranges containing samples based on the ChunkerIndex, ensuring ranges are split such that excess data is never returned.
+Determinism.
+Mixtera’s implementation of this algorithm ensures determinism because (1) the keys are processed in a consistent order and (2) when multiple component keys match a mixture key, they are considered in a deterministic order based on a seeded shuffle of all possible keys.
+This ensures that identical queries with identical mixtures always produce identical chunks, which is important for debugging and reproducibility (Zhuang et al., 2022; Cooper et al., 2022; Qian et al., 2021; Karamcheti et al., 2021; Team, 2025).
+Sampling and randomness.
+Chunks are generated randomly, but not completely i.i.d.: (i) each chunk must contain the proportions specified by the mixture , and (ii) all matching component keys are considered equal.
+Consider a mixture key that matches component keys and .
+The order in which and are processed is determined by a seeded shuffle, ensuring determinism across runs while providing randomization across different matching keys.
+Once a component key is selected, samples are drawn from its intervals.
+Within each file, intervals are used for I/O efficiency, but the order in which files are processed is shuffled.
+Importantly, if the sample count for mixture key is satisfied by samples from , samples from may never be selected for that chunk.
+This is not a sampling bias but rather reflects the mixture specification.
+By omitting the license property from , the user declares all JavaScript samples equivalent regardless of license.
+To ensure unbiased sampling across all relevant dimensions, users should include all properties they care about in the mixture.
 Mixture types.
 All mixture classes implemented in Mixtera share a common interface that converts their specifications into a mapping from MixtureKeys to sample counts per chunk, used by the chunk generation algorithm:
-Static Mixture: Users explicitly specify fixed proportions for different property combinations (Figure 4). This supports arbitrary properties and is not limited by, e.g., directory boundaries.
+Static Mixture: Users explicitly specify fixed proportions for different property combinations (Figure˜4). This supports arbitrary properties and is not limited by, e.g., directory boundaries.
 Inferring Mixture: Automatically derives mixture proportions from the data distribution in the query result: This is useful when users want to maintain the natural distribution of properties.
-Hierarchical Mixture: An advanced static mixture that allows to specify nested property relationships.
+Hierarchical Mixture: An advanced static mixture that allows specifying nested property relationships.
 For example, users can define that 50 % of the data should be legal texts, and within that, 60 % should be in English and 40% in French.
@@
 Chunk distribution.
-In distributed training, it is important to guarantee that all nodes within the same data parallel group operate on exactly the same input tensors.
+In distributed training, it is important to guarantee that all nodes within the same data parallel group operate on the same input tensors.
 Mixtera’s ChunkDistributor wraps around the chunk generation component, and hands out chunks correctly to the training nodes, i.e., the same chunks in the same order to nodes within the same group, and different chunks to nodes in different groups for data parallelism.
@@
 The protocol is message-based, with each message consisting of a task identifier followed by task-specific payload data.
-Tasks for example include the execution of a query or sending out a new chunk to a client.
+Tasks, for example, include the execution of a query or sending out a new chunk to a client.
 To handle network issues gracefully, the client implementation includes automatic reconnection with exponential backoff and configurable timeouts.
-Larger data transfers, such as chunk transmission, use efficient streaming to avoid memory issues.
+Note that only small objects such as chunks and not actual training data is transferred via Mixtera, to avoid training bottlenecks.
 5.3. Client-Side Reading
-Mixtera’s client-side abstractions offer a generator that yields data samples for a query at the server.
-This generator internally follows a two-level nested iteration pattern: an outer iteration over chunks and an inner iteration over samples within each chunk.
+Mixtera’s client-side abstractions provide a generator that, given a chunk from the server, yields the actual sample payloads .
+This generator follows a two-level nested iteration pattern: an outer iteration over chunks and an inner iteration over samples within each chunk.
 The outer iteration hides the complexity of network transfer and chunk generation, while the inner iteration hides the complexity of going from pointers in the chunk to actual samples.
-We discuss this inner step, i.e., how, given a chunk, we yield sample payloads.
+The previous section discussed the outer step, and we now discuss this inner step, i.e., how, given a chunk, we yield sample payloads.
+Sample granularity.
+A critical aspect of Mixtera’s client-side reading is understanding the sample unit.
+As defined in Section˜4.1, a sample is the atomic unit of training data as supplied by the user.
+The sample granularity is inherent to the ingested dataset .
+Mixtera is a read-only layer and does not modify or redefine this granularity.
+In practice, sample granularity varies significantly, from books to partial sentences.
+Varying granularities creates a challenge for mixture enforcement.
+If samples come in different lengths, enforcing mixture proportions at the sample level may not reflect the actual mixture at the token level seen during training.
+A single long sample from one domain may contribute orders of magnitude more tokens than a short sample from another domain.
+To address this, Mixtera provides three processing modes that offer different trade-offs between mixture guarantee granularity and sample utilization.
 Processing modes.
@@
 Tokenized mixture mode.
-This mode extends the strict window mode.
+This mode addresses the sample granularity issue identified above by enforcing mixture ratios at the token level rather than the sample level.
 It wraps the active iterators with a tokenizing iterator that takes the incoming string samples, tokenizes them, and yields tokenized samples (integer lists) with the correct sequence length.
-By setting the window size equal to the chunk size we guarantee that each chunk at least yields one window of tokenized samples.
-This mode is important for handling imbalanced collection with significantly varying text lengths.
-For example, in the The Pile (Gao et al., 2020), an average sample from the Books3 domain is much longer than one from the PubMed Abstracts domain.
-Without tokenization, one string sample may correspond to potentially order-of-magnitudes more tokenized samples, which can heavily disturb the actual mixture seen during training.
+By setting the window size equal to the chunk size we guarantee that each chunk yields at least one window of tokenized samples.
+This ensures that the mixture is respected in terms of actual training tokens, regardless of the varying lengths of the underlying samples .
 While this mode ensures precise mixture ratios at the token level, it may result in partial utilization of longer samples.
@@
 Dataset abstractions.
-Training frameworks typically require specific dataset interfaces that supports multiprocessing with worker processes.
-Besides a general purpose interface, Mixtera offers a class extending torch’s IterableDataset, and a class compatible with the huggingface API.
+Training frameworks typically require specific dataset interfaces that support multiprocessing with worker processes.
+Besides a general-purpose interface, Mixtera offers a class extending torch’s IterableDataset, and a class compatible with the Hugging Face API.
 Each worker process at each node operates on its own chunk.
@@
 Mixtera’s API offers a function to be called on checkpoint.
-In order to restore from checkpoint, we need to know (i) which chunks have been handed out to which nodes and (ii) which data loader worker processes have yielded how many samples for each node.
-Mixtera implements (i) using the ChunkDistributor, which caches the query and the current state on checkpoint and is able to restore the in-memory state of the iterators for chunk generation based on this information.
+To restore from a checkpoint, we need to know (i) which chunks have been handed out to which nodes and (ii) which data loader worker processes have yielded how many samples for each node.
+Mixtera implements (i) using the ChunkDistributor, which caches the query and the current state on checkpoint and can restore the in-memory state of the iterators for chunk generation based on this information.
 For (ii) the MixteraTorchDataset uses a shared memory segment to share with the main training process the status of the data loader workers.
@@
 […]
-We explore the first question in a dynamic mixture case study on LLMs (Section 6.1) and a static mixture case study for VLMs (Section 6.2); the second question is explored in throughput benchmarks (Section 6.3).
-We conduct our experiments on the Clariden partition of the Alps supercomputer of the Swiss National Supercomputer Centre.
-Clariden consists of HPE Cray Supercomputing EX254n blades, each hosting two Quad GH200 nodes.
+We explore the first question in a dynamic mixture case study on LLMs (Section˜6.1) and a static mixture case study for VLMs (Section˜6.2).
+We explore the second question with throughput benchmarks (Section˜6.3).
+Setup.
+We run experiments on HPE Cray Supercomputing EX254n blades, each hosting two Quad GH200 nodes.
 Each node contains 4 interconnected groups of a Grace CPU with 72 cores, 128 GB DRAM, and a H100 Hopper GPU with 96 GB of HBM.
-The nodes are connected using 200 Gb/s HPE Slingshot interconnect.
-We refer to Fusco et al.
-(Fusco et al., 2024) for a more detailed analysis of the Alps supercomputer.
-The machines run Ubuntu Server 24.01 LTS with kernel 5.14.21, and we build upon the NVIDIA NGC 25.01 container with Python 3.12, a nightly build of PyTorch 2.6, NVIDIA driver 550.54.15 and CUDA 12.8.
-We add support for Mixtera and other data loaders on our fork of torchtitan (commit d989842)222Available at https://github.com/eth-easl/torchtitan-mixtera.
+The nodes are connected using a 200 Gb/s HPE Slingshot interconnect.
+The machines run Ubuntu Server 24.04 LTS with kernel 5.14.21.
+We build on the NVIDIA NGC 25.01 container with Python 3.12, a nightly build of PyTorch 2.7, NVIDIA driver 550.54.15, and CUDA 12.8.
+We add support for Mixtera and other data loaders on our fork of Torchtitan (commit ae4e402)222Available at https://github.com/eth-easl/torchtitan-mixtera.
 (Liang et al., 2024).
 Torchtitan is part of the PyTorch ecosystem and straightforward to set up.
-We use Llama3-like models whose configurations can be found in Table 2.
+We use Llama3-like models with configurations in Table˜2.
 The 162M and 1.6B models are based on Jiang et al.
@@
 They do not have the same parameter count as torchtitan does not tie the weight embeddings.
-Our training and benchmarking data is based on The Pile (Gao et al., 2020).
-We heuristically split long samples with more than 3000 words, with max.
+Our training and benchmarking data is based on The Pile (Gao et al., 2020), a frequently-used data collection used for exploring data mixing (Jiang et al., 2024; Xie et al., 2024; Albalak et al., 2023).
+We split long samples with more than 1 500 words, with max.
 20 k samples per file.
 6.1. Dynamic Mixing using ADO
-We first demonstrate how to implement dynamic mixing algorithms in Mixtera, taking the ADO algorithm (Section 2.1.1) as an example.
-We show how dynamic mixing can improve model performance during pre-training.
-Dynamic mixing algorithms are an active area of research (Jiang et al., 2024; Chen et al., 2024a, 2023; Li et al., 2025; Albalak et al., 2023) and Mixtera facilitates exploration.
-Model Steps Mixture HellaSwag WinoGrande ARC-E ARC-C LogiQA2 Lambada (OAI) OpenBookQA PIQA SlimP.
-PPL Pile PPL 162M 15 k ADO 0.28 0.49 0.43 0.20 0.23 0.29 0.19 0.61 58.19 66.88 Default 0.28 0.52 0.44 0.19 0.23 0.29 0.18 0.61 59.13 69.52 Natural 0.28 0.52 0.41 0.18 0.23 0.26 0.16 0.61 57.54 69.85 30 k ADO 0.29 0.50 0.44 0.19 0.23 0.31 0.17 0.62 52.62 61.82 Default 0.28 0.52 0.44 0.20 0.23 0.34 0.16 0.62 53.55 64.39 Natural 0.29 0.54 0.44 0.19 0.24 0.31 0.16 0.63 52.78 65.03 1B 15 k ADO 0.37 0.54 0.51 0.23 0.25 0.52 0.20 0.70 30.07 32.20 Default 0.35 0.54 0.52 0.23 0.22 0.47 0.17 0.68 29.34 33.02 Natural 0.35 0.54 0.52 0.23 0.23 0.50 0.17 0.68 31.31 34.40 30 k ADO 0.40 0.55 0.55 0.26 0.23 0.54 0.23 0.71 25.89 29.05 Default 0.38 0.56 0.56 0.24 0.22 0.56 0.21 0.70 25.10 28.65 Natural 0.38 0.56 0.55 0.25 0.22 0.57 0.21 0.69 26.55 29.83 3B 15 k ADO 0.42 0.56 0.57 0.25 0.24 0.59 0.21 0.71 25.78 23.00 Default 0.38 0.54 0.52 0.25 0.24 0.50 0.20 0.68 26.21 24.73 30 k ADO 0.45 0.59 0.60 0.28 0.24 0.59 0.24 0.73 22.22 20.52 Default 0.42 0.56 0.59 0.26 0.24 0.57 0.23 0.70 22.52 21.79
+We demonstrate how to implement dynamic mixing algorithms in Mixtera and their impact on model performance, taking the ADO algorithm (Section˜2.1.1) as an example.
+Model Mixture HellaSwag WinoGrande ARC-E ARC-C Lambada (OAI) OpenBookQA PIQA SlimP.
+Perp.
+Pile Perp.
+1.6B ADO 0.411 0.577 0.577 0.262 0.603 0.236 0.707 25.20 27.79 Default 0.380 0.549 0.543 0.243 0.551 0.218 0.695 26.55 30.72 Natural 0.383 0.556 0.559 0.247 0.562 0.214 0.698 26.04 29.94 3.6B ADO 0.449 0.626 0.601 0.276 0.624 0.244 0.732 22.58 22.26 Default 0.415 0.597 0.579 0.255 0.592 0.210 0.706 22.09 22.02 Natural 0.419 0.586 0.593 0.276 0.598 0.206 0.723 21.99 22.53
+Training setup.
+We test the 1.6B model from Jiang et al.
+(Jiang et al., 2024) with the EleutherAI/GPT-NeoX-20B tokenizer and a 3.6B model (Table˜2) following Llama-3.2-3B from Meta, including its tokenizer.
+We omit the results from the 162M model for brevity.
+We use a sequence length of 2048.
+For ADO, we follow the codebase and discard the first 500 steps for fitting the scaling laws, start with fitting them at step 1 000, and then re-fit the laws every 1 000 steps with a subsampling frequency of 10.
+We also follow the codebase and “use the same step size” for all domains, i.e., instead of using the count of how often a domain has been sampled to fit the scaling laws like in the paper, we average the total sample counts evenly across all domains.
+We train using non-strict token-level mixtures for 30 000 steps, using a learning rate of 0.001 with a linear warmup for 500 steps and linear cooldown for 3 000 steps, and the AdamW optimizer.
+For 1.6B, we use 64 GPUs with a microbatch size of 32, and for 3.6B we use 128 GPUs with a microbatch size of 16, resulting in a global batch size of 2048 and total 125 B tokens.
+As mixtures, we test the default weights as in DoReMi (Xie et al., 2024), the natural weights as in Jiang et al.
+(Jiang et al., 2024), and ADO initialized with the natural weights.
+Evaluation metrics.
+We follow Jiang et al.
+(Jiang et al., 2024) and report both downstream task performance as well as perplexity.
+For downstream tasks, we report performance on HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-Easy and ARC-Challenge (Clark et al., 2018), Lambada OpenAI (Paperno et al., 2016), OpenBookQA (Mihaylov et al., 2018), and PIQA (Bisk et al., 2020).
+For perplexity, we report the average unweighted token perplexity on (i) the validation set of The Pile (Gao et al., 2020), and on (ii) SlimPajama (Soboleva et al., 2023) as a dataset we did not train on.
+We collect all metrics using EleutherAI’s lm-eval-harness (Gao et al., 2023), and use the unnormalized accuracy.
+ADO algorithm performance overview.
+Table˜3 shows the performance of all models and mixtures for the final checkpoint after learning rate cooldown.
+We mark in bold the best value within a model/step group.
+Generally, ADO beats the static mixtures across all downstream benchmarks.
+On 1.6B, it also has the best (lowest) perplexity on both SlimPajama and The Pile, while on 3.6B, the static mixtures have slightly lower perplexity.
+This shows that ADO leads to better performance for downstream tasks across scales, and demonstrates the usage of a dynamic mixing algorithm in Mixtera.
+Jiang et al.
+(Jiang et al., 2024) report that on the 1.6B model, ADO sometimes performs worse than static mixtures, which we do not confirm.
+Since ADO’s official repository is tightly coupled with the training framework, even after corresponding with the authors, we were not able to identify the root cause of this, partly also because their code is bound to training on specific cloud instances.
+Mixtera decouples the training framework from the mixing algorithm, which helps developers port existing algorithms to their setups (Section˜3).
+Performance over time.
+To demonstrate how different tasks behave over the training, we show the performance of the 1.6B model on HellaSwag, OpenBookQA, and ARC-Easy for all training checkpoints in Figure˜5.
+Every benchmark exhibits different behavior.
+For HellaSwag, ADO consistently increases its margin over the static mixtures.
+For OpenBookQA, ADO performs similarly to the natural mix in the intermediate checkpoints, and benefits a lot during the learning rate cooldown.
+For ARC-Easy, the mixtures perform similarly, with ADO having a small edge.
+This motivates future research on data mixing using Mixtera–for example, we might be able to use intermediate evaluations instead of loss to dynamically adjust the mixture.
+Mixture over time.
+We showcase the mixture over time for the six largest domains in Figure˜6.
+We show the mixture obtained on the 1.6B model, but unlike Jiang et al.
+(Jiang et al., 2024), we do not observe that other models/tokenizers lead to largely different mixtures.
+In all cases, the weight of GitHub and ArXiv rapidly decreases.
+Notably, the more parameters, the higher the weight of Books3, and the lower the weight of PubMed Central.
+On the 3.6B model, OpenWebText2’s weight surpasses PubMed Central’s weight before step 5 000, while for the 1.6B model, they only slowly approach.
 Mixtera implementation.
+We implement ADO in Mixtera in ca.
+800 LOC.
+The class only handles the core algorithm, while Mixtera hides the complexity of the actual mixing from the algorithm implementation.
+We implement it from scratch as the original implementation of ADO is fully tied to their training framework and data setup.
 The original ADO implementation updates the current mixture after every step and samples the next batch based on this distribution.
-The distribution is updated based on the per-domain loss from the previous step.
-This does not fully match Mixtera, which can only use a new mixture when generating a new chunk.
-Each chunk then may yield several batches of data with the same mixture.
+This is a small discrepancy to Mixtera, which can only use a new mixture when generating a new chunk.
+Each chunk may then yield several batches of data with the same mixture.
 We still send the per-domain losses on every training step at the client to update ADO’s internal state at the server.
-Whenever a new chunk is generated, the current mixture from ADO is queried (Algorithm 1), and the server generates a chunk according to that mixture.
-This introduces some slack since we do not use a new mixture on every batch.
-However, due to the stochastic nature of sampling, we do not find this to be an issue.
-The only change needed in training clients is the implementation of a per-domain loss.
+Whenever a new chunk is generated, the current mixture from ADO is queried (Algorithm˜1), and the server generates a chunk according to that mixture.
+As we find in the experiments, this slight slack does not harm performance, due to the stochastic nature of sampling.
+In order to use ADO, at the training nodes, the only change needed is the implementation of a per-domain loss.
 For this, the loss function (e.g., cross-entropy) needs to be called without reduction, which gives a loss per token.
@@
 We then perform an all-reduce operation across all training nodes to get the global per-domain losses and send this to the server.
-While the original implementation of ADO is fully tied to their training framework and data setup, Mixtera is agnostic to the training framework.
-While developing, we switched from nanotron (HuggingFace, 2025) to torchtitan (Liang et al., 2024).
-No changes in Mixtera were required.
-Only the per-domain loss module at the training framework is required.
+During development, we switched from nanotron (HuggingFace, 2025) to torchtitan (Liang et al., 2024).
+Notably, as Mixtera is agnostic to the training framework, no changes in Mixtera were required.
 This showcases the benefit of having a system like Mixtera that decouples the mixing from the training framework.
-Training setup.
-We base our setup on Jiang et al.
-(Jiang et al., 2024).
-In addition to their 162M and 1.6B architectures, to test how ADO scales to the latest model architecture, we try a 3.6B model (Table 2) following Llama-3.2-3B from Meta, including its tokenizer.
-The 162M and 1.6B models use the EleutherAI/GPT-NeoX-20B tokenizer.
-All models use a sequence length of 2048.
-For ADO, we follow the codebase and discard the first 500 steps for fitting the scaling laws, start with fitting the scaling laws at step 1 000, and then re-fit the laws every 1 000 steps with a subsampling frequency of 10.
-We also follow the codebase and “use the same step size” for all domains.
-Instead of using the count of how often a domain has been sampled to fit the scaling laws like in the paper, we follow the ADO codebase and average the total sample counts evenly across all domains.
-We train all models with non-strict token-level mixtures, a learning rate of 0.001 with a linear warmup for 500 steps and linear decay until the final step, and the AdamW optimizer, for 30 000 steps.
-For 162M/1.6B, we use 64 GPUs with a microbatch size of 32, and for 3.6B we use 128 GPUs with a microbatch size of 16, resulting in a global batch size of 2048 and total 125 B tokens.
-As mixtures, we test ADO, the default weights as in the DoReMi paper (Xie et al., 2024), and on 162M/1.6B, the “natural” weights by Jiang et al.
-(Jiang et al., 2024) which are optimized for the NeoX tokenizer.
-Evaluation metrics.
-We follow Jiang et al.
-(Jiang et al., 2024) and report both downstream task performance as well as perplexity.
-For downstream tasks, we report performance on HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-Easy and ARC-Challenge (Clark et al., 2018), LogiQA2 (Liu et al., 2023), Lambada OpenAI (Paperno et al., 2016), OpenBookQA (Mihaylov et al., 2018), and PIQA (Bisk et al., 2020).
-For perplexity, we report the average unweighted token perplexity on (i) the validation set of The Pile (Gao et al., 2020), and on (ii) SlimPajama (Soboleva et al., 2023) as a dataset we did not train on.
-We collect all metrics using EleutherAI’s lm-eval-harness (Gao et al., 2023), and use the unnormalized accuracy.
-ADO algorithm performance overview.
-Table 3 shows the performance of all models and mixtures for a checkpoint after 15 k and 30 k steps.
-We mark in bold the best value within a model/step group.
-In contrast to Jiang et al.
-(Jiang et al., 2024), we find that on the 162M model, the default mixture is quite competitive across benchmarks.
-For the final checkpoint (30 k), ADO achieves the best (lowest) perplexity on both SlimPajama and The Pile, but it does not consistently beat the static mixtures as reported in the ADO paper.
-However, analysis of the textual outputs reveals that the quality of generated text from this small model is limited across all configurations.
-For the larger 1B model, ADO is particulary strong on the intermediate checkpoint, which shows an improvement in time-to-accuracy.
-For the final checkpoint, on some benchmarks, the static mixtures perform on-par or marginally better than ADO, but on others, such as HellaSwag, ADO clearly beats the static mixtures.
-On the 3B model, ADO beats the static default mixture on both checkpoints across all benchmarks.
-While Jiang et al.
-(Jiang et al., 2024) report that ADO performs best on the smaller scale model, we find that ADO performs better on larger models.
-Since ADO’s official repository is tightly coupled with the training framework, even after corresponding with the authors, we were not able to identify the root cause of this, partly also because their code is bound to training on specific cloud instances.
-This shows that a common codebase for data mixing decoupled from the training framework is beneficial and help developers integrate the latest algorithms into their setups (Section 3).
-TODO!
-Performance over time.
-To demonstrate how different tasks behave over the training, we show the performance of the 1B model on HellaSwag, OpenBookQA, and ARC-Easy for all training checkpoints in Figure 5.
-Every benchmark exhibits different behavior.
-For HellaSwag, we can clearly see that ADO consistently increases its margin over the static mixtures.
-For OpenBookQA, ADO performs well especially in the intermediate checkpoints, and the static mixtures greatly improve towards the end.
-For ARC-Easy, all mixtures behave similarly.
-This motivates future research on data mixing using Mixtera.
-For example, we might be able to use intermediate evaluations instead of loss to dynamically adjust the mixture.
-TODO!
-Mixture over time.
-We showcase the mixture over time for the six largest domains in Figure 6.
-Around step 21 k, we run out of data for the Books3 domain, and Mixtera’s best-effort algorithm re-distributes the weight proportionally to the other domains.
-This happens because The Pile’s samples are very imbalanced in size, i.e., the average sample in Books3 has around 538 KiB, while the average Pile-CC sample has 4 KiB.
-While we split long samples heuristically, there still is some waste due to token-level mixtures.
-Unlike Jiang et al.
-(Jiang et al., 2024), we do not observe that other models/tokenizers lead to different mixtures.
-Notably, the more parameters, the higher the weight of Books3, and the lower the weight of PubMed Central.
-On the 3.6B model, OpenWebText2’s weight even surpasses PubMed Central’s weight before step 5 000.
 Takeaways.
-Dynamic mixture algorithms can improve model accuracy.
-ADO scales beyond the 1.6B model tested in the original paper, beating the default weights on the 3.6B model on all benchmarks.
+Dynamic mixtures can improve model accuracy.
+ADO scales beyond the 1.6B model tested in the original paper, and beats the default weights on all benchmarks.
 Our experiments also demonstrate that a synchronous algorithm, which uses a new mixture at each training step, adapts to Mixtera’s chunking system.
-We observe some discrepancies on the 162M model, which highlights the importance of having a common, open-source data mixing platform like Mixtera that facilities debugging and reproducibility.
 6.2. Multimodal LLaVA Finetuning
+To showcase Mixtera’s multimodal capabilities, we evaluate the impact of static mixtures for finetuning a LLaVA-style model (Liu et al., 2024).
 We are not aware of prior work on the impact of data mixtures on VLMs.
-To showcase Mixtera’s multimodal capabilities, we evaluate the impact of static mixtures for finetuning a LLaVA-style model (Liu et al., 2024).
 The LLaVA framework trains an adapter between a pre-trained image encoder and a pre-trained LLM, and then fine-tunes the adapter and LLM on visual instruction-following data.
@@
 (Liu et al., 2024; Team, 2023) and pre-train the adapter on a 558 k subset of the LAION-CC-SBU dataset with BLIP captions.
-For finetuning, we follow the TinyLLava Factory “LLaVA dataset” and use 665 k samples from six datasets (COCO (Lin et al., 2014), GQA (Hudson and Manning, 2019), OCR-VQA (Mishra et al., 2019), TextVQA (Singh et al., 2019), and VisualGenome (VG) (Krishna et al., 2017), and LLaVA’s text-only SFT annotations (Liu et al., 2024)) (Team, 2024b).
+For finetuning, we follow the TinyLLava Factory (Team, 2024b) “LLaVA dataset” and use 665 k samples from six datasets (COCO (Lin et al., 2014), GQA (Hudson and Manning, 2019), OCR-VQA (Mishra et al., 2019), TextVQA (Singh et al., 2019), and VisualGenome (VG) (Krishna et al., 2017), and LLaVA’s text-only SFT annotations (Liu et al., 2024)).
 We pre-train the adapter once, and then vary the proportions of the finetuning datasets.
@@
 Since the number of datapoints in comparison to LLM training is small, we use a best-effort mixture and ensure we go through all samples exactly once.
-All models see the same data, but in different order.
+All models see the same data, but in a different order.
+Model SQA-IMG TxtVQA GQA MME MMMU POPE Jia et al.
+(Jia et al., 2024) 64.0 49.6 58.6 1256.5 28.3 86.3 Infer.
+Mix.
+55.88 37.92 54.63 1238.76 29.4 86.6 Mix.
+#252 63.01 43.87 56.14 1268.05 30.4 85.7 Mix.
+#107 58.50 45.18 58.36 1283.62 30.1 85.54 Mix.
+#155 57.14 42.37 57.14 1290.06 29.3 86.63
 Benchmarks.
@@
 Results.
-In Table 4, we show the results reported by Jia et al.
-(Jia et al., 2024), the results we obtain using the inferring mixture (Section 5.2.2), and three mixtures out of the generated mixtures that perform well.
+In Table˜4, we show the results reported by Jia et al.
+(Jia et al., 2024), the results we obtain using the inferring mixture (Section˜5.2.2), and three mixtures out of the generated mixtures that perform well.
 While the inferring mixture does not achieve their reported results, this could be either due to different data dynamics in their training, or due to a different evaluation setup we cannot reproduce as their model weights are not public.
@@
 Overall, despite seeing the same samples globally, the mixtures play a big role for model accuracy.
-Exploring mixtures for pre-training of VLMs and adapting dynamic mixtures to VLMs are promising future work directions.
-Model SQA-IMG TxtVQA GQA MME MMMU POPE Jia et al.
-(Jia et al., 2024) 64.0 49.6 58.6 1256.5 28.3 86.3 Infer.
-Mix.
-55.88 37.92 54.63 1238.76 29.4 86.6 Mix.
-#252 63.01 43.87 56.14 1268.05 30.4 85.7 Mix.
-#107 58.50 45.18 58.36 1283.62 30.1 85.54 Mix.
-#155 57.14 42.37 57.14 1290.06 29.3 86.63
 6.3. Throughput Benchmarks
 We now focus on throughput and compare Mixtera with other data loaders across various configurations.
-Mixtera’s goal is to avoid data stalls, i.e., training throughput should not be reduced because we are waiting for data (Mohan et al., 2021; Murray et al., 2021; Kuchnik et al., 2022; Zhao et al., 2022).
-Hence, we measure throughput in tokens per second in actual workloads.
-Note that smaller models and more data parallelism increase the pressure on the data loader, while larger models reduce the pressure.
+The goal is avoiding data stalls, i.e., training throughput should not be reduced because the GPU is waiting for data (Mohan et al., 2021; Kuchnik et al., 2022; Murray et al., 2021; Zhao et al., 2022; Graur et al., 2022; Böther et al., 2025; Robroek et al., 2026; Chen et al., 2025).
+We want to show that Mixtera enables on-the-fly data streaming with dynamic mixing without lowering throughput in comparison to other data loaders.
+We measure throughput in tokens per second.
+Note that smaller models and more data parallelism increase the pressure on the data loader, while larger models reduce the pressure as the training computation takes longer.
 If a data loader can sustain training small models at scale on a high-end platform like the GH200, its performance is sufficient for other scenarios as well.
-Besides Mixtera with chunk sizes of 512, 1024, and 2048, we benchmark the iterable HuggingFaceDataset by Torchtitan (Hf-Iter), a mapped version (Hf-Map), the Mosaic StreamingDataset (Team, 2022) (Mosaic).
+We run Mixtera with chunk sizes of 512, 1024, and 2048.
+We compare to well-known state-of-the-art data loaders, i.e., the iterable HuggingFaceDataset by Torchtitan (Hf-Iter), a mapped version (Hf-Map), and the Mosaic StreamingDataset (Team, 2022) (Mosaic).
 The difference between Hf-Iter and Hf-Map is that similar to Mixtera, Hf-Iter loads and tokenizes the data on the fly, while Hf-Map preprocesses all data, including tokenization.
-We evaluate throughput on the 162M model since larger models only lead to lower throughput.
+Note that none of these data loaders support dynamic mixing (Table˜1), they just read files from front to back.
+We evaluate throughput on the 162M model since larger models only lead to lower throughput, as discussed above.
 We always use FSDP since Torchtitan only enables bfloat16 training if sharding is used.
@@
 We use the Hugging Face EleutherAI/gpt-neox-20b tokenizer for all data loaders.
-We store all data on the Iopsstor SSD-backed Lustre DFS in the Alps cluster.
-TODO!
+We store all data on an SSD-backed Lustre DFS.
 Single-node.
@@
 We use 2 data parallel replicates and shards.
-In Figure 7 we show the throughput for the data loaders depending on the number of data workers.
+In Figure˜7 we show the throughput for the data loaders depending on the number of data workers, i.e., data loading processes from the torch.DataLoader.
 We test up to 16 workers, where 0 workers indicate that data is loaded in the same process as the main training loop.
-We only show results up to 4 workers due as throughput does not increase further.
-The error bars show the standard deviation of the throughput.
+The plot only show results up to 4 workers as throughput does not increase further.
 Without data workers, Mixtera has the highest average throughput; with one worker, the other data loaders reach their peak performance.
-Mixtera performs worse than the other loaders with a lower number of workers due to the random access into the files.
+Overall, Mixtera provides similar throughput to the baselines while having a much richer feature set, e.g., dynamic mixing.
+Notably, Mixtera has higher throughput variance than other loaders for lower number of workers.
+This is due to the random access into the files.
 For every sample, it needs to open the file, seek to the correct position, and load the data, instead of bulk-transferring all the data as the other data loaders can do.
-This leads to the higher variance in throughput indicated by the error bars, which overall leads to a lower average.
-However, with a higher number of workers, the variance in data access gets hidden.
-When using 4+ workers, Mixtera performs almost identical to the other data loaders.
+This leads to the higher variance in throughput indicated by the error bars, which overall leads to slightly lower averages.
+With a higher number of workers and larger chunk sizes, the variance decreases.
+When using 4+ workers, Mixtera’s throughput matches the throughput of the baselines.
 Additionally, increasing the chunk size also helps, and for the 0 worker case, Mixtera with a chunk size of 1024 or 2048 even has a higher average throughput.
 This benchmark setup is quite extreme, as we train a very small 162M model on an extremely fast GPU.
-Nevertheless, just by adding some workers, which comes at no overhead beyond some CPU resources, Mixtera reaches competitive throughput despite its more complicated model.
+For larger, state-of-the-art model sizes, the throughput differences disappear, as computation time spent within the model forward and backward passes increases.
+Mixtera’s client-side overhead is minimal; the primary operations (chunk parsing and data fetching) introduce negligible overhead.
 Scaling out.
@@
 For 4, 8, and 16 GPUs, we use half the GPUs as replication, and shard across the rest.
-The results with 8 data workers can be found in Figure 8.
-All data loaders scale perfectly linear.
-For larger number of GPUs, the throughput variance increases a bit for all systems.
+Figure˜8 shows the results with 8 data workers.
+All data loaders scale linearly.
+With more GPUs, the throughput variance increases a bit for all systems.
 We attribute this to the random assignment of nodes in the cluster by the Slurm scheduler across the 3 repetitions.
 We do not test pipeline or tensor parallelism as (i) this is not necessary for the 162M model (and a larger model would only stress the data loaders less), (ii) the data loaders besides Mixtera do not easily support 3D parallelism, (iii) increasing data parallelism increases the load on the system more.
+This demonstrates the scalability of Mixtera’s single-controller design and implementation, as well as the efficiency of its chunk generation, enabling it to effectively supply chunks to all clients even at scale.
 File formats.
@@
 This observation holds across different numbers of data workers, where we also observe minimal performance variation among the data loaders.
-TODO!
-6.3.1. Query execution
-Mixtera executes the query in DuckDB and generates the ChunkerIndex before we can start training.
-On our Pile dataset with 241 M samples, DuckDB takes 29 s, and preparing the index takes 18 s due to our C++ implementation.
-Optionally, we can also persist an initial state checkpoint of the query, which due to serialization of nested Python dictionaries takes 13 min.
-This is only necessary if checkpointing should be supported.
-Future checkpoints can then be stored in milliseconds.
-While the streaming Hf-Iter data loader can basically start streaming data immediately, the Hf-Map data loader, the default in nanotron, loads and tokenizes the data first, taking 2 h 51 min for this dataset.
-However, what needs to be considered is that for streaming data loaders such as Hf-Iter or Mosaic, in many scenarios, users would also need to run more offline preprocessing, e.g., to perform static filtering, or reshuffling the data if we want to mix on a different property.
-Mixtera avoids this offline preprocessing completely, and, besides an optional state checkpoint, starts streaming data in under a minute.
+Predicates.
+The previous benchmarks select all domains within The Pile.
+To show how a filtering predicate impacts throughput, in Figure˜9, we show the throughput for chunks sizes 512/2048, and 0/2 data workers, across four representative predicates: Books3 only (large samples), CommonCrawl (CC) and ArXiv (frequent domains with small and large samples), CC only (frequent but small samples), and a set of infrequent domains (Enron Emails, NIH ExPorter, PhilPapers, EuroParl, USPTO).
+With 0 data workers, throughput varies with the predicate; more workers stabilize throughput.
+For chunk size 512, Only CC achieves ~587 kTokens/s, while Books3 reaches ~776 kTokens/s.
+These differences stem from different file access patterns.
+Because The Pile is shuffled, large domains are more likely to appear in clusters, reducing repeated file scans.
+Domains also differ in sample sizes: large text samples yield multiple training examples from a single read, lowering total data transfer, as seen with Books3.
+With two workers, Mixtera masks these effects and throughput is uniform.
+Clustering.
+If data is clustered, Mixtera can generate longer intervals and thereby reduce the amount of data transfers.
+Figure˜10 compares the throughput of the regular benchmarking data with a clustered version that sequentially writes out each domain’s data.
+For a small number of workers, we generally observe an increase in avg.
+throughput and a slight decrease of variance.
+As before, using more workers hides these effects.
+6.4. Data Ingestion
+Before executing queries, the user must ingest metadata (c.f.
+Section˜5.1).
+During ingestion, Mixtera reads all samples, extracts their metadata, and inserts it into the underlying metadata database.
+This cost is paid only once; thereafter, arbitrarily many queries can be executed.
+In Figure˜11 we report time spent reading the samples (scanning all files) and inserting metadata.
+We benchmark this for varying dataset sizes (10 %, 100 %, and 250 % of The Pile) and number of attributes (randomly generated to avoid correlation).
+Other operations (e.g., parser setup) have negligible overhead.
+The timings are obtained using Mixtera’s optimized default settings: multithreaded file scanning (MT) and an insertion chunk size of 2,000.
+Disabling MT increases the ingestion time from 315 to around 3,000 seconds for the standard 100 % / 1 attribute case.
+With MT enabled, decreasing the chunk size to 500 increases ingestion time from 315 s to 340 s.
+Chunk sizes larger than 2,000 yield only marginal benefit and increase the risk of out-of-memory errors.
+6.5. Query Execution Latency Breakdown
+When a query is executed in Mixtera, the main steps are querying the populated ODC (DuckDB), building the ChunkerIndex, and, optionally, writing the first checkpoint.
+In Figure˜12, we give a time breakdown for varying dataset sizes (10 %, 100 %, and 250 % of The Pile) and number of properties (randomly generated to avoid correlation), following Section˜6.4.
+For the 100 % / 1 attribute case, DuckDB takes 32 s, and preparing the index takes 16 s due to our C++ implementation.
+Persisting the initial state checkpoint of the query dominates the runtime with 282 s.
+This is because we have to serialize nested Python dictionaries, which is slow despite engineering optimizations.
+Importantly, after this initial checkpoint, future checkpoints at the server can be stored within milliseconds, since we do not have to serialize the index again.
+If no Mixtera checkpoints are needed, the serialization can be skipped.
+Other systems.
+The streaming Hf-Iter data loader can basically start streaming data immediately, the Hf-Map data loader, the default in nanotron, loads and tokenizes the data first, taking 2 h 51 min for The Pile 100 %.
+However, to use streaming data loaders such as Hf-Iter or Mosaic, in many scenarios, users would also need to run more offline preprocessing, e.g., to perform static filtering, or reshuffling the data if we want to mix on a different property.
+Mixtera avoids this offline preprocessing completely.
 7. Conclusion and Future Work
-Managing training data at scale requires data management support.
-We present Mixtera, a training framework-agnostic data plane for foundation model training.
-Using its chunking mechanism, it supports dynamic mixtures across arbitrary properties.
-We demonstrate Mixtera’s performance and scalability, as well as the importance of mixtures on both LLMs and VLMs.
-For future work, it is interesting to extend the analysis to the cloud, i.e., using a service like S3 as the underlying storage layer.
-Mixtera also lays the foundation to implement lineage tracking on which model was trained on which data, as it is a centralized access point.
-From a ML perspective, the impact of data composition and quality on model performance is still under-explored, and Mixtera provides a foundation to conduct future research on dynamic and static mixing algorithms.
+Understanding how data mixture recipes affect model quality is an active and critical area of ML research.
+We design Mixtera to enable researchers and model developers to easily and quickly train models with a variety of mixtures across arbitrary data properties and dynamically vary mixtures during training.
+Mixtera is a declarative data plane for foundation model training that is training framework-agnostic.
+We demonstrate Mixtera’s throughput and scalability, as well as the impact of mixtures on model quality for both LLMs and VLMs.
+The system lays the foundation for implementing additional features, such as data lineage tracking for model training, as it has a global view of the data, and enables future research on compute-intensive modalities such as video (Ye et al., 2025) and audio (Zhao et al., 2025).
 Acknowledgements.
 We thank Beste Güney for her contributions to Mixtera’s codebase.
-We thank Antoni-Joan Solergibert i Llaquet, Imanol Schlag, Steven Hand, Martin Jaggi, Antoine Bosselut, Alexander Hägele, Loubna Ben Allal, Quentin de Laroussilhe, Paul Barham, Yiding Jiang, Theodoros Rekatsinas, and Foteini Strati for helpful discussions.
+We thank Antoni-Joan Solergibert i Llaquet, Imanol Schlag, Steven Hand, Martin Jaggi, Antoine Bosselut, Alexander Hägele, Loubna Ben Allal, Quentin de Laroussilhe, Paul Barham, Yiding Jiang, Theodoros Rekatsinas, Gustavo Alonso, and Foteini Strati for helpful discussions.
 This work was supported as part of the Swiss AI Initiative by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID a09 on Alps.
@@
 doi:10.48550/arXiv.2502.02737
-- Armbrust et al.
-(2021) Michael Armbrust, Ali Ghodsi, Reynold Xin, and Matei Zaharia.
-2021.
-Lakehouse: A New Generation of Open Platforms that Unify Data Warehousing and Advanced Analytics.
-In Proceedings of the Conference on Innovative Data Systems Research (CIDR).
+- Barham et al.
+(2022) Paul Barham, Aakanksha Chowdhery, Jeff Dean, Sanjay Ghemawat, Steven Hand, Dan Hurt, Michael Isard, Hyeontaek Lim, Ruoming Pang, Sudip Roy, Brennan Saeta, Parker Schuh, Ryan Sepassi, Laurent El Shafey, Chandramohan A.
+Thekkath, and Yonghui Wu.
+2022.
+Pathways: Asynchronous Distributed Dataflow for ML.
+In Proceedings of the Conference on Machine Learning and Systems (MLSys).
 - Ben-Nun and Hoefler (2019) Tal Ben-Nun and Torsten Hoefler.
@@
 doi:10.1609/aaai.v34i05.6239
+- Bodensohn et al.
+(2025) Jan-Micha Bodensohn, Ulf Brackmann, Liane Vogel, Anupam Sanghi, and Carsten Binnig.
+2025.
+Unveiling Challenges for LLMs in Enterprise Data Engineering.
+arXiv preprint (2025).
+doi:10.48550/arXiv.2504.10950
 - Bommasani et al.
@@
 doi:10.48550/arXiv.2108.07258
+- Böther et al.
+(2025) Maximilian Böther, Ties Robroek, Viktor Gsteiger, Xianzhe Ma, Pınar Tözün, and Ana Klimovic.
+2025.
+Modyn: Data-Centric Machine Learning Pipeline Orchestration.
+In Proceedings of the Conference on Management of Data (SIGMOD).
+doi:10.1145/3709705
 - Brown et al.
@@
 doi:10.48550/arXiv.2307.14430
+- Chen et al.
+(2025) Zihao Chen, Chenyang Zhang, Chen Xu, Zhao Zhang, Jiaqiang Wang, Weining Qian, and Aoying Zhou.
+2025.
+Scheduling Data Processing Pipelines for Incremental Training on MLP-based Recommendation Models.
+In Companion of the International Conference on Management of Data (SIGMOD).
+doi:10.1145/3722212.3724454
 - Clark et al.
@@
 doi:10.1145/3511265.3550446
-- ETH AI Center & EPFL AI Center (2025) ETH AI Center & EPFL AI Center. 2025. Swiss AI Initiative. https://www.swiss-ai.org/.
 - Fu et al.
@@
 doi:10.48550/ARXIV.2306.13394
-- Fusco et al.
-(2024) Luigi Fusco, Mikhail Khalilov, Marcin Chrapek, Giridhar Chukkapalli, Thomas C.
-Schulthess, and Torsten Hoefler.
-2024.
-Understanding Data Movement in Tightly Coupled Heterogeneous Systems: A Case Study with the Grace Hopper Superchip.
-arXiv preprint (2024).
-doi:10.48550/ARXIV.2408.11556
 - Gao et al.
@@
 doi:10.5281/zenodo.10256836
+- Graur et al.
+(2022) Dan Graur, Damien Aymon, Dan Kluser, Tanguy Albrici, Chandramohan A Thekkath, and Ana Klimovic.
+2022.
+Cachew: Machine Learning Input Data Processing as a Service.
+In Proceedings of the USENIX Annual Technical Conference (ATC).
 - Hoefler et al.
@@
 doi:10.1109/cvpr52733.2024.02484
-- Liu et al.
-(2023) Hanmeng Liu, Jian Liu, Leyang Cui, Zhiyang Teng, Nan Duan, Ming Zhou, and Yue Zhang.
-2023.
-LogiQA 2.0—An Improved Dataset for Logical Reasoning in Natural Language Understanding.
-IEEE/ACM Transactions on Audio, Speech, and Language Processing 31 (2023).
-doi:10.1109/taslp.2023.3293046
 - Longpre et al.
@@
 - Richburg and Carpuat (2024) Aquia Richburg and Marine Carpuat. 2024. How Multilingual Are Large Language Models Fine-Tuned for Translation? arXiv preprint (2024). doi:10.48550/arXiv.2405.20512
+- Robroek et al.
+(2026) Ties Robroek, Neil Kim Nielsen, and Pınar Tözün.
+2026.
+TensorSocket: Shared Data Loading for Deep Learning Training.
+In Proceedings of the Conference on Management of Data (SIGMOD).
+doi:10.1145/3749185
 - Sakaguchi et al.
@@
 - Ye et al.
+(2025) Juncheol Ye, Seungkook Lee, Hwijoon Lim, Jihyuk Lee, Uitaek Hong, Youngjin Kwon, and Dongsu Han.
+2025.
+SAND: A New Programming Abstraction for Video-based Deep Learning.
+In Proceedings of the Symposium on Operating Systems Principles (SOSP).
+doi:10.1145/3731569.3764847
+- Ye et al.
 (2024) Jiasheng Ye, Peiju Liu, Tianxiang Sun, Yunhua Zhou, Jun Zhan, and Xipeng Qiu.
@@
 - Zhao et al.
+(2025) Juntao Zhao, Qi Lu, Wei Jia, Borui Wan, Lei Zuo, Junda Feng, Jianyu Jiang, Yangrui Chen, Shuaishuai Cao, Jialing He, Kaihua Jiang, Yuanzhe Hu, Shibiao Nong, Yanghua Peng, Haibin Lin, Xin Liu, and Chuan Wu.
+2025.
+OVERLORD: Ultimate Scaling of DataLoader for Multi-Source Large Foundation Model Training.
+arXiv preprint (2025).
+doi:10.48550/arXiv.2504.09844
+- Zhao et al.
 (2022) Mark Zhao, Niket Agarwal, Aarti Basant, Buğra Gedik, Satadru Pan, Mustafa Ozdal, Rakesh Komuravelli, Jerry Pan, Tianshu Bao, Haowei Lu, Sundaram Narayanan, Jack Langman, Kevin Wilfong, Harsha Rastogi, Carole-Jean Wu, Christos Kozyrakis, and Parik Pol.